### v1

# Identifying and Solving Conditional Image Leakage in Image-to-Video Diffusion Model

Abstract
Diffusion models have obtained substantial progress in image-to-video (I2V) generation. However, such models are not fully understood. In this paper, we report a significant but previously overlooked issue in I2V diffusion models (I2V-DMs), namely, conditional image leakage. I2V-DMs tend to over-rely on the conditional image at large time steps, neglecting the crucial task of predicting the clean video from noisy inputs, which results in videos lacking dynamic and vivid motion. We further address this challenge from both inference and training aspects by presenting plug-and-play strategies accordingly. First, we introduce a training-free inference strategy that starts the generation process from an earlier time step to avoid the unreliable late-time steps of I2V-DMs, as well as an initial noise distribution with optimal analytic expressions (Analytic-Init) by minimizing the KL divergence between it and the actual marginal distribution to effectively bridge the training-inference gap. Second, to mitigate conditional image leakage during training, we design a time-dependent noise distribution for the conditional image, which favors high noise levels at large time steps to sufficiently interfere with the conditional image. We validate these strategies on various I2V-DMs using our collected open-domain image benchmark and the UCF101 dataset. Extensive results demonstrate that our methods outperform baselines by producing videos with more dynamic and natural motion without compromising image alignment and temporal consistency. The project page: https://cond-image-leak.github.io/.
1 Introduction
Image-to-video (I2V) generation aims to generate videos with dynamic and natural motion while maintaining the appearance of the given image. I2V generation allows users to guide video creation from the input image (and optional text), thus increasing controllability and flexibility in content creation. Like the remarkable progress in text-to-image (T2I) generation  [40, 38, 20, 8, 18, 5] and text-to-video (T2V) generation  [10, 42, 6, 11], diffusion models have also obtained promising results for I2V generation [9, 15, 54, 12, 58]. However, such models are not fully understood.
Figure 1: Schematic diagram of conditional image leakage. Ideally, I2V-DMs predict clean videos from noisy inputs, using the conditional image as auxiliary content guidance. However, at large time steps, the heavily corrupted input retains minimal video detail, causing the model tends to over-rely on the detailed conditional image and neglect the crucial task of synthesizing video from noisy inputs.
In this paper, we identify a significant but previously overlooked issue in the image-to-video diffusion models (I2V-DMs), namely, conditional image leakage (see Sec. 3.1). Ideally, the I2V-DMs should predict the added noise or clean video mainly from noisy inputs, with the conditional image serving as auxiliary guidance. However, as the diffusion process progresses—especially at large time steps, the heavily corrupted input retains minimal video detail, while the conditional image preserves extensive detail of the target video. This biases the model to over-rely on the conditional image while neglecting the essential task of synthesizing video from noisy inputs, leading to videos lacking dynamic and vivid motion. For instance, I2V-DMs that take a shortcut by directly replicating the conditional image may even achieve a lower loss. To validate this issue, we corrupt a ground truth (GT) clean video data using the forward transition kernel and compute one-step clean video predictions across all time steps. As illustrated in Fig. 2, predicted clean videos exhibit markedly reduced motion than GT at large time steps, with videos becoming almost entirely static at the final step.
Based on the analysis, we attempt to address the challenge in both inference and training. First, we present a simple yet effective training-free inference strategy, where we start the video generation process from an earlier time step instead of the final one, thus avoiding the unreliable late-time steps of the I2V-DMs (see Sec. 4.1). To further enhance performance, we derive an initial noise distribution with optimal analytic expressions (Analytic-Init) by minimizing the KL divergence between it and the actual marginal distribution to bridge the training-inference gap. Second, to mitigate the conditional image leakage during training, we introduce a time-dependent noise distribution (TimeNoise) that interferes with the conditional image, particularly at large time steps (Sec. 4.2). In principle, the distribution should favor high noise levels at large time steps to sufficiently disrupt the conditional image, then gradually decrease the noise levels as the time step diminishes. To achieve this, we utilize a logit-normal distribution with a center that monotonically increases over time and conduct a systematic analysis of its two hyperparameters. Notably, our strategies are designed to be plug-and-play, making them adaptable to various I2V-DMs based on both VP-SDE [12, 54] and VE-SDE [9] frameworks. Finally, the clean videos predicted by our method maintain motion dynamics comparable to the GT across all time steps, indicating the absence of conditional image leakage (see Fig. 6).
Empirically, we validate our methods on various I2V-DMs [9, 54, 12] using our collected open-domain images (ImageBench) and the UCF101 dataset. We conduct a user study on the ImageBench and report FVD [49] and IS [7] on UCF101. Extensive experimental results demonstrate that our strategies outperform baselines by producing videos with markedly more dynamic and natural motion, all while maintaining image alignment and temporal consistency.
2 Background
Diffusion Models.
Diffusion models gradually perturb data ${\bm{x}}_{0}\sim q({\bm{x}}_{0})$ through a forward diffusion process and reserve it to recover the data. The forward transitional kernel $q_{t|0}({\bm{x}}_{t}|{\bm{x}}_{0})$ is given by
${\bm{x}}_{t}=\alpha_{t}{\bm{x}}_{0}+\sigma_{t}{\bm{\epsilon}},\quad{\bm{\epsilon}}\sim{\mathcal{N}}({\bm{0}},{\bm{I}}),\quad t\in[0,T],$ (1)
where $\alpha_{t}$ and $\sigma_{t}$ are the noise schedule and chosen to ensure that ${\bm{x}}_{T}$ contains minimal information about ${\bm{x}}_{0}$. Two prevalent types of forward stochastic differential equation (SDEs) are commonly used [44, 28]. One is the variance-preserving SDE (VP-SDE) [40, 23], where $\alpha_{t}^{2}+\sigma_{t}^{2}=1$ with $\alpha_{t}\rightarrow 0$ as $t\rightarrow T$, ensuring $p_{T}({\bm{x}}_{T})=\mathcal{N}({\bm{0}},{\bm{I}})$. The other is the variance exploding SDE (VE-SDE), where $\alpha_{t}=1$ and $\sigma_{T}$ is set to a large constant, resulting in $p_{T}({\bm{x}}_{T})={\mathcal{N}}({\bm{0}},\sigma_{T}^{2}{\bm{I}})$. Such models can be parameterized using a noise-prediction model $\epsilon_{\theta}({\bm{x}}_{t},t)$ (${\bm{\epsilon}}$-prediction) [23] to optimize:
$\mathbb{E}_{{\bm{x}}_{0}\sim q({\bm{x}}_{0}),{\bm{\epsilon}}\sim{\mathcal{N}}({\bm{0}},{\bm{I}}),t\sim\mathcal{U}(1,T)}\left[\|{\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t)-{\bm{\epsilon}}\|_{2}^{2}\right],$ (2)
where the noisy input ${\bm{x}}_{t}\sim q_{t|0}({\bm{x}}_{t}|{\bm{x}}_{0})$. Alternative parametrizations such as ${\bm{x}}_{0}$-prediction [28], ${\bm{v}}$-prediction [41], $F$-prediction [27] are also commonly used. The ${\bm{\epsilon}}$-prediction and ${\bm{x}}_{0}$-prediction loss aims to predict the added noise or clean video from noisy input ${\bm{x}}_{t}$. Starting from ${\bm{x}}_{T}\sim p_{T}({\bm{x}}_{T})$, various samplers [43, 32, 33, 27] can be employed to generate data.
Diffusion Models for Image-to-Video Generation.
Given any image ${\bm{y}}_{0}$ from the open domain, the goal of I2V is to generate a video $X_{0}=\{{\bm{x}}_{0}^{i}\}_{i=1}^{N}$ with dynamic and natural motion while maintaining the appearance of ${\bm{y}}_{0}$. This task can be formulated as designing a conditional distribution $p_{\theta}(X_{0}|{\bm{y}}_{0})$, which be achieved by a conditional diffusion model:
$\mathbb{E}_{X_{0},{\bm{y}}_{0},{\bm{\epsilon}},t}\left[\|{\bm{\epsilon}}_{\theta}(X_{t},{\bm{y}}_{0},t)-{\bm{\epsilon}}\|_{2}^{2}\right],$ (3)
where $X_{t}\sim q_{t|0}(X_{t}|X_{0})$. Typically, ${\bm{y}}_{0}$ is the first frame of $X_{0}$ and DynamiCrafter [54] adopts a randomly selected frame from $X_{0}$ as ${\bm{y}}_{0}$. The key issue is effectively integrating the conditional image ${\bm{y}}_{0}$ into the diffusion model. Most methods use CLIP image embeddings [39] to maintain the semantic content of ${\bm{y}}_{0}$. Notably, VideoCrafter1 [12] and Dynamicrafter [54] employ the last layer’s full patch visual tokens from the CLIP ViT, enriching the encoded information, and other approaches prefer the class token layer. Yet, solely depending on these embeddings, such as in VideoCrafter1 [12], compromises detail retention, impacting image alignment. To enhance detail representation, I2VGen-XL [58] combines the conditional image with the noisy initial frame, while VideoComposer [51] develops a STC-encoder for multiple conditions. Although superior to CLIP image embeddings, these strategies still fail to fully retain the conditional image content. To mitigate this, AnimateAnything [15], Dynamicrafter [54] and SVD [9] choose to directly concatenate noisy video $X_{t}$ with ${\bm{y}}_{0}$, which provides detailed information.
Figure 2: Identifying conditional image leakage in I2V-DMs. We corrupt a ground truth (GT) clean video and use it to compute $\hat{X}_{t\rightarrow 0}$ prediction across time steps. As time progresses, $\hat{X}_{t\rightarrow 0}$ exhibits markedly reduced motion than GT, with the videos becoming nearly static at the final step $T$, suggesting the existence of conditional image leakage in a large time step.
3 Conditional Image Leakage
In this section, we first identify a commonly overlooked issue in I2V-DMs: conditional image leakage, discussed in Sec. 3.1. We then discuss existing I2V-DMs with respect to this issue in Sec. 3.2.
3.1 Identifying Conditional Image Leakage in Image-to-video Diffusion Models
Although existing I2V-DMs discussed in Sec. 2 have obtained significant progress, such models are not fully understood. This section revisits and underscores a critical yet previously overlooked issue: conditional image leakage (CIL). As shown in Fig. 1, ideally, I2V-DMs should predict the added noise ($\epsilon$-prediction) or clean video (${\bm{x}}_{0}$-prediction) mainly from noisy input $X_{t}$, using the conditional image ${\bm{y}}_{0}$ as auxiliary content guidance. However, as the diffusion process progresses—particularly at large time steps, the heavily corrupted input $X_{t}$ contains minimal information about $X_{0}$, making it challenging to predict clean frames. Meanwhile, the conditional image retains considerable detail of $X_{0}$. This biases the model to over-rely on the conditional image while neglecting the essential task of synthesizing video from noisy inputs. For instance, a diffusion model that takes a shortcut by directly replicating ${\bm{y}}_{0}$ can even achieve a lower loss than one that accurately predicts the clean video from a heavily corrupted $X_{t}$. As a result, the synthesized videos lack dynamic and vivid motion. Notably, recent strategies that involve adjusting the noise schedule towards more noise [14, 26, 9, 31] may further exacerbate this issue (see details in Appendix C).
To validate this issue, we corrupt a ground truth (GT) clean video $X_{0}$ using the forward transition kernel in Eq. (1) and use it as the noisy input to compute the one-step prediction $\hat{X}_{t\rightarrow 0}$ at time $t$:
$\hat{X}_{t\rightarrow 0}=\left(X_{t}-\sigma_{t}{\bm{\epsilon}}_{\theta}(X_{t},y,t)\right)/\alpha_{t}.$ (4)
Ideally, $\hat{X}_{t\rightarrow 0}$ attempts to predict the GT $X_{0}$ from noisy input $X_{t}$ and exhibit comparable motion dynamics. However, as shown in Fig. 2, as time progresses—particularly at large time steps, $\hat{X}_{t\rightarrow 0}$ exhibits markedly reduced motion than GT, with the videos becoming nearly static at the final step $T$, suggesting the existence of the conditional image leakage. Consequently, it will generate an almost still video starting from $T$ by duplicating the conditional image (see Fig. 6).
Figure 3: Understanding exiting work from conditional image leakage. (a) I2VGen-XL [58] mitigates the leakage at the expense of image alignment. (b) Increasing the SVD motion score induces unnatural camera movements with stationary objects. Our inference strategy, however, produces dynamic and natural motion in both people and the teapot.
3.2 Understanding Existing Work from Conditional Image Leakage
We now examine three popular I2V-DMs with respect to conditional image leakage: I2VGEN-XL [58], Dynamicrafter [54], and SVD [9]. Although these models do not explicitly address conditional image leakage, we believe their strategies may mitigate it to some extent.
Firstly, I2VGEN-XL [58] adds ${\bm{y}}_{0}$ with the first frame of noisy input, which reduces information about ${\bm{y}}_{0}$ largely and thus mitigates the leakage issue. However, Fig. 3(a) indicates that the generated videos do not accurately preserve ${\bm{y}}_{0}$. Dynamicrafter [54] and SVD [9] directly feed ${\bm{y}}_{0}$ into I2V-DMs, enhancing detail preservation but also facilitating leakage. Dynamicrafter selectively refines spatial layers while preserving the pre-trained temporal layers, which contain motion priors and thus maintain motion dynamics to a certain extent. However, it does not inherently solve the leakage issue. SVD [9] introduces an additional motion score to compel the model to generate videos with motion to align with motion score, thereby reducing reliance solely on ${\bm{y}}_{0}$ and mitigating conditional image leakage. However, we argue that it does not address the fundamental challenge facing I2V-DMs, where I2V-DMs should generate clean video primarily from noisy input. An example in Fig. 3(b) highlights this limitation. By increasing the motion score, the method produces videos with significant camera movement to meet the scalar motion score but leaves objects static. In contrast, the strategy we propose in Sec. 4.1 results in a video with vivid and naturalistic movement, where both the people and the teapot exhibit dynamic motion, offering a more effective solution to the fundamental challenges of I2V-DMs. Furthermore, if an inappropriate motion score is provided during inference, the resulting video may exhibit unnatural motion or poor temporal consistency to match the given score (see Fig. 12 in the Appendix). In summary, despite efforts by these works to somewhat mitigate conditional image leakage, the issue persists, underscoring the ongoing need for further exploration.
Figure 4: Benefits of tuning $M$ and Analytic-Init w.r.t. visual quality. An appropriate $M$ can enhance performance by increasing motion without compromising other performance. A too-small $M$ delivers poor visual quality due to the training-inference gap. Analytic-Init helps to mitigate it.
4 Solving Conditional Image Leakage in Image-to-video Diffusion Models
In this section, we tackle the issue of conditional image leakage from two perspectives: first, by outlining a training-free inference strategy to mitigate it, and second, by presenting a training strategy designed to rectify this phenomenon during the training phase.
4.1 Inference Strategy
As discussed in Sec. 3.1, conditional image leakage commonly occurs at large time steps. To address it, a straightforward solution is to start the generation process from an earlier time step $M\in[0,T]$, avoiding the unreliable later stages of I2V-DMs. Let $p_{M}(X_{M})$ denote the initial noise distribution at the start time $M$. Initially, we set $p_{M}(X_{M})=p_{T}(X_{T})$, i.e. ${\mathcal{N}}({\bm{0}},{\bm{I}})$ in VP-SDE [23] or ${\mathcal{N}}({\bm{0}},\sigma_{T}^{2}{\bm{I}})$ in VE-SDE [28]. As illustrated in the third column ($M=0.92T$) of Fig. 4, this straightforward strategy markedly improves motion dynamics without sacrificing other performance. However, a smaller $M$ value (e.g., $M=0.8T$) hurts the visual quality due to the training-inference discrepancy.
To bridge the training-inference gap, we propose to minimize the KL divergence between the initial noise distribution $p_{M}(X_{M})$ and the actual marginal distribution of the diffusion process at time $M$, denoted as $q_{M}(X_{M})$. Inspired by existing work [4, 3, 55], when $p_{M}(X_{M})$ is modeled as a normal distribution ${\mathcal{N}}(X_{M};{\bm{\mu}}_{p},\sigma_{p}^{2}{\bm{I}})$, we derive analytical expressions for the optimal mean ${\bm{\mu}}_{p}^{*}$ and variance $\sigma_{p}^{2\ast}$ as follows. We refer to it as the Analytic Noise Initialization (Analytic-Init).
Proposition 1.
Given a normal distribution $p_{M}(X_{M})={\mathcal{N}}(X_{M};{\bm{\mu}}_{p},\sigma_{p}^{2}{\bm{I}})$ and $q_{M}(X_{M})$ is the margin distribution of diffusion forward process at time $M$, with the forward trainsition kernel $q_{M|0}(X_{M}|X_{0})={\mathcal{N}}(X_{M};\alpha_{M}X_{0},\sigma_{M}^{2}{\bm{I}})$, the minimization problem $\underset{{\bm{\mu}}_{p},\sigma_{p}^{2}}{\min}D_{KL}(q_{M}(X_{M})||p_{M}(X_{M}))$ yields the following optimal solution:
${\bm{\mu}}_{p}^{*}=\alpha_{M}\mathbb{E}_{q(X_{0})}[X_{0}],\quad\sigma_{p}^{2\ast}=\alpha_{M}^{2}\frac{\sum_{j=1}^{d}[Var(X_{0}^{(j)})]}{d}+\sigma_{M}^{2},$ (5)
where $q(X_{0})$ denotes the data distribution , $d$ denotes the dimension of the data, and $X_{0}^{(j)}$ denotes the j-th component of $X_{0}$.
Appendix A provides detailed proof of Proposition 1. Empirically, we can estimate ${\bm{\mu}}_{p}^{*}$ and $\sigma_{p}^{2\ast}$ using the method of moments [4]. Compared with the naive approach where $p_{M}(X_{M})=p_{T}(X_{T})$, Analytic-Init can achieve superior visual quality by reducing the training-inference gaps, especially with a smaller $M$. The theoretical insights and Proposition 1 are validated by the results in Tab. 4, measured by the commonly used FVD and IS metrics.
Figure 5: Effects of tuning $\mu(t)$ and $\beta_{m}$ in TimeNoise. Adjusting these parameters allows for a flexible trade-off between dynamic motion and image alignment. Higher noise levels result in more pronounced motion but reduced image alignment.
Figure 6: Effectiveness of TimeNoise to address conditional image leakage. We replicate the experiments described in Sec. 3.1 to evaluate TimeNoise. Applying TimeNoise ensures that $\hat{X}_{t\rightarrow 0}$ maintains motion dynamics comparable to the GT across all time steps, indicating the absence of conditional image leakage. Consequently, we can generate videos with high motion starting from $T$.
4.2 Training Strategy
The previous section explores strategies to mitigate conditional image leakage during inference. This section presents methods to address this issue during the training phase. As noted in Sec. 3.1, the conditional image ${\bm{y}}_{0}$ remains informative of the target video, leading I2V-DMs to depend excessively on it. Therefore, an intuitive idea is to perturb ${\bm{y}}_{0}$ to relieve the reliance on it.
Our first attempt is to introduce noise similar in scale to that used in $X_{t}$. This aims to equalize the model’s challenge in predicting clean video from $X_{t}$ or ${\bm{y}}_{t}$, reducing its dependency on ${\bm{y}}_{0}$. However, this strategy also makes it difficult to harness ${\bm{y}}_{0}$, resulting in lower video quality. To overcome this, we propose a noise distribution on ${\bm{y}}_{0}$ that introduces substantial noise to prevent leakage while maintaining a cleaner ${\bm{y}}_{0}$ to aid content generation. As discussed earlier, as time progresses, $X_{t}$ becomes less informative about $X_{0}$, increasing the risk of leakage. This insight leads us to develop a time-dependent noise distribution $p_{t}(\beta_{s})$ (TimeNoise). The design principle is to favor high noise levels at large time steps to sufficiently disrupt ${\bm{y}}_{0}$, shifting towards lower noise levels as the time step decreases. To achieve this, we employ a logit-normal distribution [18, 1] defined as below:
$p_{t}(\beta_{s};\mu(t),\beta_{m})=\frac{\beta_{m}}{\sqrt{2\pi}}\frac{1}{\beta_{s}(\beta_{m}-\beta_{s})}e^{-\frac{(\text{logit}(\frac{\beta_{s}}{\beta_{m}})-\mu(t))^{2}}{2}},$ (6)
where $\text{logit}(\frac{\beta_{s}}{\beta_{m}})$ follows a normal distribution centered around $\mu(t)$ with a standard deviation of 1. Our noise distribution includes two hyperparameters: $\beta_{m}$, the maximum noise level, and $\mu(t)$, the center of the distribution. As illustrated in Fig. 8(a), we can adhere to the previously stated design principle of $p_{t}(\beta_{s})$ by adjusting $\mu(t)$ over time $t$. Finally, the noisy conditional image ${\bm{y}}_{s}$ at time $t$ is obtained by ${\bm{y}}_{s}={\bm{y}}_{0}+\beta_{s}{\bm{\epsilon}}$, where $\beta_{s}\sim p_{t}(\beta_{s}),{\bm{\epsilon}}\sim{\mathcal{N}}({\bm{0}},{\bm{I}})$. We also tried other adding noise choices but found them to be less effective (see Appendix D).
Next, we conduct a systematic analysis of two hyperparameters to investigate their impact on video generation. The first, $\mu(t)$, should ideally increase monotonically with time, where we set it ranging from $\mu(0)=-1$ to $\mu(T)=1$. We explore three typical functional forms for $\mu(t)$ (see Fig. 8(b)): (1) linear: $\mu(t)=2t-1$, (2) convex:$\mu(t)=2t^{a}-1$, where $a>1$ and (3) concave: $\mu(t)=2t^{a}-1$, where $a\in(0,1)$. As shown in Fig. 5(a), our TimeNoise strategy produces videos with more dynamic motion than the baseline. Functions that introduce higher noise levels, such as the concave function, enhance dynamic motion but reduce temporal consistency and image alignment. We also test the following baselines: (1) constant, where $\beta_{s}=\beta_{m}\frac{\mu(t)+1}{2}$ represents a single value rather than a distribution, and (2) a conditioning augmentation method from CDM [24]. Results indicate that the former method results in poor image alignment, while the latter shows low motion (see Fig. 10 in the Appendix). We then examine the maximum noise level, $\beta_{m}$. As shown in Fig. 5(b), a higher $\beta_{m}$ enhances dynamic motion but decreases temporal consistency and image alignment, while a lower $\beta_{m}$ leads to less motion. Considering that VideoCrafter1 and DynamiCrafter utilize full patch visual tokens from CLIP, not just the class token, we also apply noise to these tokens. Finally, we replicated the experiments outlined in Sec. 3.1, and the results depicted in Fig. 6 demonstrate that our TimeNoise ensures $\hat{X}_{t\rightarrow 0}$ maintains motion dynamics comparable to the ground truth across all time steps. This indicates an absence of conditional image leakage, underscoring the effectiveness of our approach.
Figure 7: Visualizing the effectiveness of our inference and training strategies. Ours significantly enhances video dynamism while maintaining image alignment and temporal consistency. VC. and DC. denote VideoCrafter1 [12] and DynamiCrafter [54] respectively.
5 Related Work
Diffusion Models for Image Generation.
Recently, diffusion models have recently achieved significant breakthroughs in image, video, and 3D generation [23, 16, 40, 37, 5, 59, 36, 52, 53, 60]. For image generation, the latent diffusion model [40] addresses computational costs by leveraging VQ-VAE [47] to transform pixel-space images into compact latent representations, subsequently training diffusion models within this latent space. Building upon the latent space concept, subsequent works such as SDXL [38], DALLE-3 [8], and SD3 [18] have further enhanced performance. Exploiting these advancements in text-to-image diffusion models, numerous methods have demonstrated promising results in text-driven controllable image generation and image editing [57, 21, 46]. In image editing, studies such as Prompt-to-Prompt [21] and Plug-and-Play [46] have explored attention-based control mechanisms over generated content, consistently delivering impressive results. For image translation, EGSDE [59] and DiffuseIT [29] propose to employ an additional energy function to guide the inference process of a pre-trained diffusion model.
Diffusion Models for T2V Generation.
Approaches to T2V generation can be classified into two main categories. The first category involves directly generating videos based on text [42, 10, 22, 12, 13, 48, 56, 50, 9, 17]. For instance, as a pioneering endeavor, Make-A-Video [42] utilizes a pre-trained text-to-image model along with a prior network for T2V diffusion models, obviating the necessity for paired video-text datasets. VideoLDM [10] maintains the parameters of a pre-trained T2I model while fine-tuning newly introduced temporal layers. More recently, transformers have emerged as a foundational architecture for video generation due to their scalability [20, 35, 34]. The second category typically entails a two-step generation process: first generating an image based on the textual input and subsequently creating a video conditioned on the text and the generated image [30, 19]. For instance, Emu-video [19] initializes a factorized text-to-video model using a pre-trained text-to-image model and then fine-tunes temporal modules in the I2V stage.
[…]
Figure 8: (a)Visualization of designed noise distribution. Adjusting $\mu(t)$ over time $t$ can satisfy the design principle. For example, set large $\mu(t)$ for large $t$ (purple curve) and small $\mu(t)$ for small $t$ (blue curve). (b)Visualization of different $\mu(t)$.
6 Experiments
Datasets and Evaluation Metrics.
We validate our training strategy on the WebVid-2M dataset [2], resizing and center-cropping all videos to $320\times 512$ with 16 frames. For evaluation, we validate our methods on UCF101 and our collected ImageBench, comprising 100 images from various websites and T2I models like SDXL [38] and UniDiffuser [5], covering categories like nature, humans, and more for comprehensive evaluation. For ImageBench, we conduct user studies with 10 subjects to perform pairwise comparisons of our methods against baselines, evaluating motion, temporal consistency, image alignment, and overall performance. On UCF101, we report Fréchet Video Distance (FVD) and Inception Score (IS). Further details can be found in Appendix B.
Implementation Details.
For Analytic-Init, we use 5000 samples from the Webvid-2M dataset to estimate $p_{M}(X_{M})$ by default. We set $M=0.96T,0.96T$ and $\sigma_{M}=100$ for VideoCrafter1 [12], DynamiCrafter [54], and SVD [9] on UCF101. For ImageBench, we adjust $M$ to $0.92T,0.92T$ and $\sigma_{M}=100$. For the TimeNoise, we set $\beta_{m}=25,100$, and $100$ for VideoCrafter1, DynamiCrafter, and SVD, respectively. The function $\mu(t)=2t^{5}-1$ is applied across all baselines. These baselines are fine-tuned for 20,000 iterations, using either the official code or a replication of the official settings except for batch size. More detailed information can be found in the Appendix B.
[…]
6.1 Results
The effectiveness of our inference and training strategy.
We validate our strategies on three I2V-DMs: VideoCrafter1 [12], DynamiCrafter [54], and SVD [48]. Demonstrated in Fig. 7, our strategies significantly enhance video dynamism and natural motion, all while maintaining image alignment and temporal consistency. The findings are corroborated by the user study in Fig. 9. Considering overall aspects, ours achieves superior results. Tab. 8 showcases the quantitative superiority of our inference and training strategies, which outperform all baselines significantly. For instance, on DynamiCrafter [54], we improve the FVD score by 47.5 and 47.6 using our inference and training strategy respectively, suggesting the effectiveness of our method.
Evaluating the combined inference and training strategies.
In this section, we aim to investigate the necessity of combining two strategies. As illustrated in Tab. 3, for DynamiCrafter [54] and SVD [9], our TimeNoise effectively mitigates conditional image leakage, rendering the additional inference strategy less impactful. Conversely, for VideoCrafter1 [12], which relies solely on CLIP image embedding for information, employing excessive TimeNoise disrupts image alignment. Hence, we utilize a moderate TimeNoise ($\beta_{m}=25$), where the inference strategy remains effective.
7 Conclusions and Discussions
In this paper, we identify a common issue in I2V-DMs: conditional image leakage. We address this challenge from two aspects. First, we introduce a training-free inference strategy that starts the generation process from an earlier time step to avoid the unreliable late-time steps of I2V-DMs. Second, we design a time-dependent noise distribution for the conditional image to mitigate conditional image leakage during training. We validate the effectiveness of these strategies across various I2V-DMs.
Limitations and broader impact.
One limitation of this paper is the need to balance time-dependent noise distribution to prevent conditional image leakage while maintaining image integrity. While we demonstrate the effectiveness of our training strategy on existing image-to-video diffusion models, we do not provide a definitive noise distribution choice for a scratch-trained model. We leave this in the future work. Furthermore, we must use the method responsibly to prevent any negative social impacts, such as the creation of misleading fake videos.
Appendix A Proof of Proposition 1
According to Lemma 2 of Analytic-DPM [4], the KL divergence between initial noise distribution $p_{M}(X_{M})$ and the actual marginal distribution $q_{M}(X_{M})$ can be expressed as:
$D_{KL}(q_{M}(X_{M})||p_{M}(X_{M}))=D_{KL}({\mathcal{N}}(X_{M};{\bm{\mu}}_{q},\Sigma_{q})||p_{M}(X_{M}))$
$+H({\mathcal{N}}(X_{M};{\bm{\mu}}_{q},\Sigma_{q}))-H(q_{M}(X_{M})),$ (7)
where ${\bm{\mu}}_{q},\Sigma_{q}$ denote expectation and covariance matrix of $q_{M}(X_{M})$, and $H(\cdot)$ denotes the entropy of a distribution. Since $q_{M}(X_{M})$ is independent of $p_{M}(X_{M})$, the last two terms can be considered as constants. Therefore, minimizing the Eq. (A) is equivalent to minimizing $D_{KL}(N(X_{M};{\bm{\mu}}_{q},\Sigma_{q})||p_{M}(X_{M}))$. For normal distributions, $D_{KL}(N(X_{M};{\bm{\mu}}_{q},\Sigma_{q})||p_{M}(X_{M}))$ could be expanded as:
$\frac{1}{2}[({\bm{\mu}}_{q}-{\bm{\mu}}_{p})^{T}\Sigma_{p}^{-1}({\bm{\mu}}_{q}-{\bm{\mu}}_{p})-\log\det(\Sigma_{p}^{-1}\Sigma_{q})+tr(\Sigma_{p}^{-1}\Sigma_{q})-d],$ (8)
$=\frac{1}{2}[\frac{1}{\sigma_{p}^{2}}\lVert{\bm{\mu}}_{p}-{\bm{\mu}}_{q}\rVert^{2}+d\log(\sigma_{p}^{2})+\frac{1}{\sigma_{p}^{2}}tr(\Sigma_{q})-\log\det(\Sigma_{q})-d],$ (9)
where $d$ denotes the dimension of the data. Minimizing Eq. (9) yields:
${\bm{\mu}}_{p}^{*}={\bm{\mu}}_{q}=\mathbb{E}_{q_{M}(X_{M})}[X_{M}].$ (10)
Taking the derivative of Eq. 9 with respect to $\sigma_{p}^{2}$, we can obtain
$\sigma_{p}^{2\ast}$ $=\frac{trCov_{q_{M}(X_{M})[X_{M}]+\lVert{\bm{\mu}}_{p}-{\bm{\mu}}_{q}\rVert^{2}}}{d},$ (11)
$=\frac{trCov_{q_{M}(X_{M})[X_{M}]}}{d},$ (12)
$=\frac{\sum_{j=1}^{d}(Var(X_{M}^{(j)}))}{d}.$ (13)
Taking the $X_{M}$ defined in Eq. (1) into Eq. (10), the optimal ${\bm{\mu}}_{p}^{*}$ could be further represented as
$\mu_{p}^{*}$ $=\alpha_{M}\mathbb{E}_{q(X_{0})}[X_{0}]+\mathbb{E}_{q({\bm{\epsilon}})}[{\bm{\epsilon}}],$ (14)
$=\alpha_{M}\mathbb{E}_{q(X_{0})}[X_{0}].$ (15)
Similarly, for the optimal variance $\sigma_{p}^{2\ast}$, we can derive
$Var(X_{M}^{(j)})=Var(\alpha_{M}X_{0}^{(j)}+\sigma_{M}{\bm{\epsilon}}^{(j)})$ (16)
Given that $X_{0}$ and ${\bm{\epsilon}}$ are independent of each other, the variance $Var(X_{M}^{(j)})$ can be decomposed further:
$Var(X_{M}^{(j)})$ $=Var(\alpha_{M}X_{0}^{(j)})+Var(\sigma_{M}{\bm{\epsilon}}^{(j)}),$ (17)
$=\alpha_{M}^{2}Var(X_{0}^{(j)})+\sigma_{M}^{2}.$ (18)
Finally, taking Eq. (18) into Eq. (13), the optimal $\sigma_{p}^{2\ast}$ can be represented as
$\sigma_{p}^{2\ast}=\alpha_{M}^{2}\frac{\sum_{j=1}^{d}[Var(X_{0}^{(j)})]}{d}+\sigma_{M}^{2}.$ (19)
Appendix B Implementation Details
Table 1: Code Links and Licenses.
Method Link License
VideoCrafter1 [12] https://github.com/AILab-CVC/VideoCrafter Apache License
DynamiCrafter [54] https://github.com/Doubiiu/DynamiCrafter Apache License
SVD https://github.com/Stability-AI/generative-models MIT License
UniDiffusers [5] https://github.com/thu-ml/unidiffuser AGPL-3.0 license
SDXL https://github.com/Stability-AI/generative-models Open RAIL++-M
B.1 Code Used and License
We validate our strategies on DynamiCrafter [54] ($320\times 512$ version), VideoCrafter1 [12] ($320\times 512$ version), and SVD [9]. All used codes in this paper and their licenses are listed in Tab. 1.
B.2 Training and Inference Details
We utilize the official training code of DynamiCrafter (refer to Tab. 1) to fine-tune both DynamiCrafter [54] and VideoCrafter1 [12], and reproduce the training code for SVD by ourselves. Throughout the training phase, we maintained consistent settings across all models, with the sole exception of incorporating our TimeNoise component for a fair comparison. Each model was fine-tuned using the WebVid-2M dataset [2], where videos were resized and center-cropped to dimensions of $320\times 512$ and segmented into sequences of 16 frames. In light of the discussions in Sec. 3.2 regarding the adverse effects of motion scores and the lack of a precise method by SVD [9] to compute these scores, we set a fixed motion score of $20$ during training. We fix the frame rate at $3$ fps and use dynamic frame rates for DynamiCrafter [54] and VideoCrafter1 [12]. Additional details can be found in Tab. B.2 and Tab. B.2. Our experiments were conducted using A800-80G GPUs, and the computational costs are detailed in Tab. 2. For inference, the official model codes were used for sampling (see Tab. 1). Specifically, we employed a DDIM sampler with 50 steps for DynamiCrafter [54] and VideoCrafter1 [12], and Heun’s 2nd order method with 25 steps for SVD [9].
[…]
tableTraining settings for DynamiCrafter [54] and VideoCrafter1 [12]. Config Value Optimizer AdamW Learning rate 1e-5 Weight decay 1e-2 Optimizer momentum $\beta_{1},\beta_{2}$=0.9, 0.999 Batch size 64 Training iterations 20,000
[…]
tableTraining settings for SVD [9]. Config Value Optimizer AdamW Learning rate 3e-5 Weight decay 1e-2 Optimizer momentum $\beta_{1},\beta_{2}$=0.9,0.999 Batch size 48 Training iterations 20,000
Table 2: Compute resources.
Model Iterations GPU-type GPU-nums Hours
DynamiCrafter [54] 20,000 A800 8 8
VideoCrafter1 [12] 20,000 A800 8 8
SVD [9] 20,000 A800 6 7
B.3 Evaluation
FVD and IS. Following prior studies [25, 10], we compute the Fréchet Video Distance (FVD) and Inception Score (IS) for 2,048 and 10,000 samples on the UCF101 dataset, respectively. Specifically, the FVD is calculated using the code available at https://github.com/SongweiGe/TATS/ with a pre-trained I3D action classification model, which can be accessed at https://www.dropbox.com/scl/fi/c5nfs6c422nlpj880jbmh/i3d_torchscript.pt?rlkey=x5xcjsrz0818i4qxyoglp5bb8&dl=1. The IS is derived using the code from https://github.com/pfnet-research/tgan2, employing a pre-trained C3D model [45]. For this process, we sample 16 frames at 3 fps, resize them to the default resolution of each model, and use the first frame as the conditional image to generate videos. The FVD and IS are then computed between the generated videos and the ground truth videos. For DynamiCrafter [54] and VideoCrafter1 [12], which utilize text as an additional condition, we employ categories as the textual input.
User study. We ask users to compare ours with the baselines and determine which ones exhibit more dynamic and natural motion, greater temporal consistency, better alignment with the conditional image, and overall preference.
Table 3: Evaluating the combined inference (Inference.) and training (Train.) strategies. Rat. denotes a user study that ranks methods. Refer to Sec. 6.1 for detailed analysis.
Inference. Train. VideoCrafter1 [12] DynamiCrafter [54] SVD [9]
[…]
460.3 23.98 4 382.5 21.12 4 311.0 22.03 4
450.1 24.50 1.9 342.9 22.71 2.9 277.1 22.18 2.9
452.2 24.62 2.7 334.9 21.42 1.4 272.2 23.01 1.6
443.7 25.11 1.4 332.1 22.84 1.7 272.4 25.18 1.5
Appendix C The Influence of Adjusting the Noise Schedule
In this section, we first examine commonly used strategies that involve adjusting the noise schedule to higher noise levels to bridge the training-inference gap [14, 26, 9, 31], which, unfortunately, may exacerbate conditional image leakage. Following the SVD [9], we increase noise by adjusting the $P_{mean}$ in the EDM framework [27]. Surprisingly, as shown in Fig. 11, we observe that larger $P_{mean}$ values correspond to reduced motion in the generated videos. We hypothesize that this is due to the increased noise added to $X_{t}$ by larger $P_{mean}$, making it more challenging to predict clean frames and thus more prone to conditional image leakage. Consequently, the synthesized videos tend to lack dynamic and vivid motion.
Appendix D Ablation Studies for Adding Noise Method
In this section, we show another choice to add noise to the conditional image. Specifically, the noisy conditional image ${\bm{y}}_{s}$ at time $t$ is obtained by
${\bm{y}}_{s}=(1-\beta_{s}){\bm{y}}_{0}+\beta_{s}{\bm{\epsilon}},$ (20)
where $\beta_{s}\sim p_{t}(\beta_{s}),{\bm{\epsilon}}\sim{\mathcal{N}}({\bm{0}},{\bm{I}})$, and $\beta_{m}=1$. As shown in Fig. 13, this choice leads to video discoloration in SVD [9].
Figure 10: The qualitative comparison between our TimeNoise and baselines mentioned in Sec. 4.2. The constant results in poor image alignment, while the CDM [24] shows low motion. Ours achieves the best visual quality.
Figure 11: Adjusting the noise schedule towards more noise further exacerbate conditional image leakage.
Figure 12: An inappropriate motion score leads to poor temporal consistency.
Figure 13: Other choice of adding noise on conditional image. Adding noise via Eq. 20 causes video discoloration in SVD [9].
Table 4: Effects of tuning $M$ and Analytic-Init.
[…]
DynamiCrafter [54] T 363.8 16.39
0.98T 345.1 16.57
0.96T 343.1 17.54
0.94T 325.2 19.12
0.92T 365.8 20.49
0.9T 442.2 22.81
DynamiCrafter + Analytic-Init 0.98T 324.3 16.44
0.96T 316.3 17.66
0.94T 319.6 19.13
0.92T 347.2 20.58
0.9T 378.6 22.40
Table 5: Quantitative results of the inference strategy across varying initial time $M$ on the UCF101 dataset.
[…]
DynamiCrafter [54] T 363.8 16.39
+ Analytic-Init 0.96T 316.3 17.66
0.92T 347.2 20.58
0.88T 407.2 23.32
0.84T 535.1 24.27
0.8T 696.1 23.67
VideoCrafter1 [12] T 353.9 18.75
+ Analytic-Init 0.96T 341.6 19.86
0.92T 344.3 21.58
0.88T 368.4 21.82
0.84T 400.8 21.90
0.8T 445.6 21.21
[…]
SVD-finetune [9] 700 311.0 22.03
+ Analytic-Init 500 301.9 22.00
300 290.2 21.94
100 277.1 22.18
70 272.5 22.27
50 295.5 21.89

### v2

@@
 Abstract
-Diffusion models have obtained substantial progress in image-to-video (I2V) generation.
-However, such models are not fully understood.
-In this paper, we report a significant but previously overlooked issue in I2V diffusion models (I2V-DMs), namely, conditional image leakage.
-I2V-DMs tend to over-rely on the conditional image at large time steps, neglecting the crucial task of predicting the clean video from noisy inputs, which results in videos lacking dynamic and vivid motion.
-We further address this challenge from both inference and training aspects by presenting plug-and-play strategies accordingly.
-First, we introduce a training-free inference strategy that starts the generation process from an earlier time step to avoid the unreliable late-time steps of I2V-DMs, as well as an initial noise distribution with optimal analytic expressions (Analytic-Init) by minimizing the KL divergence between it and the actual marginal distribution to effectively bridge the training-inference gap.
-Second, to mitigate conditional image leakage during training, we design a time-dependent noise distribution for the conditional image, which favors high noise levels at large time steps to sufficiently interfere with the conditional image.
-We validate these strategies on various I2V-DMs using our collected open-domain image benchmark and the UCF101 dataset.
-Extensive results demonstrate that our methods outperform baselines by producing videos with more dynamic and natural motion without compromising image alignment and temporal consistency.
+Diffusion models have obtained substantial progress in image-to-video generation.
+However, in this paper, we find that these models tend to generate videos with less motion than expected.
+We attribute this to the issue called conditional image leakage, where the image-to-video diffusion models (I2V-DMs) tend to over-rely on the conditional image at large time steps.
+We further address this challenge from both inference and training aspects.
+First, we propose to start the generation process from an earlier time step to avoid the unreliable large-time steps of I2V-DMs, as well as an initial noise distribution with optimal analytic expressions (Analytic-Init) by minimizing the KL divergence between it and the actual marginal distribution to bridge the training-inference gap.
+Second, we design a time-dependent noise distribution (TimeNoise) for the conditional image during training, applying higher noise levels at larger time steps to disrupt it and reduce the model’s dependency on it.
+We validate these general strategies on various I2V-DMs on our collected open-domain image benchmark and the UCF101 dataset.
+Extensive results show that our methods outperform baselines by producing higher motion scores with lower errors while maintaining image alignment and temporal consistency, thereby yielding superior overall performance and enabling more accurate motion control.
 The project page: https://cond-image-leak.github.io/.
 1 Introduction
-Image-to-video (I2V) generation aims to generate videos with dynamic and natural motion while maintaining the appearance of the given image.
-I2V generation allows users to guide video creation from the input image (and optional text), thus increasing controllability and flexibility in content creation.
-Like the remarkable progress in text-to-image (T2I) generation  [40, 38, 20, 8, 18, 5] and text-to-video (T2V) generation  [10, 42, 6, 11], diffusion models have also obtained promising results for I2V generation [9, 15, 54, 12, 58].
+Image-to-video (I2V) generation aims to generate videos with dynamic and natural motion while maintaining the content of the given image.
+It allows users to guide video creation from the input image (and optional text), thus increasing controllability and flexibility in content creation.
+Like the remarkable progress in text-to-image (T2I) generation  [42, 39, 21, 8, 18, 5] and text-to-video (T2V) generation  [10, 46, 6, 11], diffusion models have also obtained promising results for I2V generation [9, 15, 63, 12, 68, 69].
 However, such models are not fully understood.
-Figure 1: Schematic diagram of conditional image leakage.
-Ideally, I2V-DMs predict clean videos from noisy inputs, using the conditional image as auxiliary content guidance.
-However, at large time steps, the heavily corrupted input retains minimal video detail, causing the model tends to over-rely on the detailed conditional image and neglect the crucial task of synthesizing video from noisy inputs.
-In this paper, we identify a significant but previously overlooked issue in the image-to-video diffusion models (I2V-DMs), namely, conditional image leakage (see Sec.
+Figure 1: The issue of existing I2V-DMs.
+Regardless of input motion scores (Input MS), the output motion scores (Output MS) are consistently lower than expected.
+In contrast, our method yields output motion scores either higher or lower than Input MS with reduced error.
+In this paper, we observe that existing image-to-video diffusion models (I2V-DMs) tend to generate videos with less motion than expected (see Fig.
+1).
+We attribute this to a previously overlooked issue called conditional image leakage (see Sec.
 3.1).
-Ideally, the I2V-DMs should predict the added noise or clean video mainly from noisy inputs, with the conditional image serving as auxiliary guidance.
-However, as the diffusion process progresses—especially at large time steps, the heavily corrupted input retains minimal video detail, while the conditional image preserves extensive detail of the target video.
-This biases the model to over-rely on the conditional image while neglecting the essential task of synthesizing video from noisy inputs, leading to videos lacking dynamic and vivid motion.
-For instance, I2V-DMs that take a shortcut by directly replicating the conditional image may even achieve a lower loss.
-To validate this issue, we corrupt a ground truth (GT) clean video data using the forward transition kernel and compute one-step clean video predictions across all time steps.
-As illustrated in Fig.
-2, predicted clean videos exhibit markedly reduced motion than GT at large time steps, with videos becoming almost entirely static at the final step.
-Based on the analysis, we attempt to address the challenge in both inference and training.
-First, we present a simple yet effective training-free inference strategy, where we start the video generation process from an earlier time step instead of the final one, thus avoiding the unreliable late-time steps of the I2V-DMs (see Sec.
-4.1).
-To further enhance performance, we derive an initial noise distribution with optimal analytic expressions (Analytic-Init) by minimizing the KL divergence between it and the actual marginal distribution to bridge the training-inference gap.
-Second, to mitigate the conditional image leakage during training, we introduce a time-dependent noise distribution (TimeNoise) that interferes with the conditional image, particularly at large time steps (Sec.
-4.2).
-In principle, the distribution should favor high noise levels at large time steps to sufficiently disrupt the conditional image, then gradually decrease the noise levels as the time step diminishes.
-To achieve this, we utilize a logit-normal distribution with a center that monotonically increases over time and conduct a systematic analysis of its two hyperparameters.
-Notably, our strategies are designed to be plug-and-play, making them adaptable to various I2V-DMs based on both VP-SDE [12, 54] and VE-SDE [9] frameworks.
-Finally, the clean videos predicted by our method maintain motion dynamics comparable to the GT across all time steps, indicating the absence of conditional image leakage (see Fig.
-6).
-Empirically, we validate our methods on various I2V-DMs [9, 54, 12] using our collected open-domain images (ImageBench) and the UCF101 dataset.
-We conduct a user study on the ImageBench and report FVD [49] and IS [7] on UCF101.
-Extensive experimental results demonstrate that our strategies outperform baselines by producing videos with markedly more dynamic and natural motion, all while maintaining image alignment and temporal consistency.
+Normally, the noisy input contains the motion information of the target video, and I2V-DMs should rely on it to predict motion, while the static conditional image provides content guidance.
+However, in practice, as the diffusion process progresses—especially at large time steps, the noisy input becomes heavily corrupted, while the conditional image preserves extensive detail of the target video.
+This biases the model to over-rely on the conditional image and neglect the noisy input, leading to videos with reduced motion.
+To validate this, we corrupt the ground truth (GT) clean video and compute one-step clean video predictions at each time step.
+As shown in Fig.
+2, the predicted clean videos exhibit markedly reduced motion than GT at large time steps, indicating leakage.
+Based on the above analysis, we attempt to address the challenge from both inference and training aspects (see Sec.
+3.2).
+First, we present a simple yet effective inference strategy that starts the video generation process from an earlier time step, thus avoiding the unreliable large-time steps of the I2V-DMs.
+To further enhance performance, we derive an initial noise distribution with optimal analytic expressions (Analytic-Init) by minimizing the KL divergence between it and the true marginal distribution to bridge the training-inference gap.
+Second, to mitigate leakage during training, we propose a time-dependent noise distribution (TimeNoise) that increases noise levels at larger time steps, effectively disrupting the conditional image and reducing model dependency on it.
+This is achieved by employing a logit-normal distribution with a center that gradually shifts over time.
+Finally, our method achieves higher motion scores with reduced motion score error and ensures that the predicted clean video maintains motion dynamics comparable to the ground truth across all time steps, effectively mitigating conditional image leakage.
+Notably, our general strategies are adaptable to various I2V-DMs based on both VP-SDE [12, 63, 69] and VE-SDE [9] framework.
+Empirically, we validate our methods on various I2V-DMs [9, 63, 12, 69] using our collected open-domain images (ImageBench) and the UCF101 dataset.
+We conduct a user study on ImageBench and report FVD [55], IS [7], and motion score [9] on UCF101.
+For motion-conditioned models [9, 69], we also report the motion score error between the generated video and the input motion score at different levels.
+Extensive experimental results demonstrate that our strategies outperform baselines by producing higher motion scores with lower errors while maintaining image alignment and temporal consistency, thereby yielding superior overall performance and enabling more accurate motion control.
 2 Background
 Diffusion Models.
-Diffusion models gradually perturb data ${\bm{x}}_{0}\sim q({\bm{x}}_{0})$ through a forward diffusion process and reserve it to recover the data.
+Diffusion models gradually perturb the data ${\bm{x}}_{0}\sim q({\bm{x}}_{0})$ via a forward diffusion process and reverse the process to recover it.
 The forward transitional kernel $q_{t|0}({\bm{x}}_{t}|{\bm{x}}_{0})$ is given by
 ${\bm{x}}_{t}=\alpha_{t}{\bm{x}}_{0}+\sigma_{t}{\bm{\epsilon}},\quad{\bm{\epsilon}}\sim{\mathcal{N}}({\bm{0}},{\bm{I}}),\quad t\in[0,T],$ (1)
-where $\alpha_{t}$ and $\sigma_{t}$ are the noise schedule and chosen to ensure that ${\bm{x}}_{T}$ contains minimal information about ${\bm{x}}_{0}$.
-Two prevalent types of forward stochastic differential equation (SDEs) are commonly used [44, 28].
-One is the variance-preserving SDE (VP-SDE) [40, 23], where $\alpha_{t}^{2}+\sigma_{t}^{2}=1$ with $\alpha_{t}\rightarrow 0$ as $t\rightarrow T$, ensuring $p_{T}({\bm{x}}_{T})=\mathcal{N}({\bm{0}},{\bm{I}})$.
-The other is the variance exploding SDE (VE-SDE), where $\alpha_{t}=1$ and $\sigma_{T}$ is set to a large constant, resulting in $p_{T}({\bm{x}}_{T})={\mathcal{N}}({\bm{0}},\sigma_{T}^{2}{\bm{I}})$.
-Such models can be parameterized using a noise-prediction model $\epsilon_{\theta}({\bm{x}}_{t},t)$ (${\bm{\epsilon}}$-prediction) [23] to optimize:
+where $\alpha_{t}$ and $\sigma_{t}$ are the noise schedule chosen to ensure that ${\bm{x}}_{T}$ contains minimal information about ${\bm{x}}_{0}$.
+Such forward diffusion processes can be viewed as stochastic differential equations (SDEs), among which two prevalent types are commonly used [48, 29].
+One is the variance-preserving SDE (VP-SDE) [42, 24], where $\alpha_{t}^{2}+\sigma_{t}^{2}=1$ with $\alpha_{t}\rightarrow 0$ as $t\rightarrow T$, ensuring $p_{T}({\bm{x}}_{T})=\mathcal{N}({\bm{x}}_{T};{\bm{0}},{\bm{I}})$.
+The other is the variance exploding SDE (VE-SDE), where $\alpha_{t}\equiv 1$ and $\sigma_{T}$ is set to a large constant, resulting in $p_{T}({\bm{x}}_{T})\approx{\mathcal{N}}({\bm{x}}_{T};{\bm{0}},\sigma_{T}^{2}{\bm{I}})$.
+Such models can be parameterized with a noise-prediction model $\epsilon_{\theta}({\bm{x}}_{t},t)$ (${\bm{\epsilon}}$-prediction) [24],and the parameters are learned by minimizing:
 $\mathbb{E}_{{\bm{x}}_{0}\sim q({\bm{x}}_{0}),{\bm{\epsilon}}\sim{\mathcal{N}}({\bm{0}},{\bm{I}}),t\sim\mathcal{U}(1,T)}\left[\|{\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t)-{\bm{\epsilon}}\|_{2}^{2}\right],$ (2)
 where the noisy input ${\bm{x}}_{t}\sim q_{t|0}({\bm{x}}_{t}|{\bm{x}}_{0})$.
-Alternative parametrizations such as ${\bm{x}}_{0}$-prediction [28], ${\bm{v}}$-prediction [41], $F$-prediction [27] are also commonly used.
-The ${\bm{\epsilon}}$-prediction and ${\bm{x}}_{0}$-prediction loss aims to predict the added noise or clean video from noisy input ${\bm{x}}_{t}$.
+Alternative parametrizations such as ${\bm{x}}_{0}$-prediction [29], ${\bm{v}}$-prediction [45], $F$-prediction [28] are also commonly applied.
+The ${\bm{\epsilon}}$-prediction and ${\bm{x}}_{0}$-prediction aims to predict the added noise or clean video from noisy input ${\bm{x}}_{t}$.
 Starting from ${\bm{x}}_{T}\sim p_{T}({\bm{x}}_{T})$, various samplers [47, 33, 34, 28] can be employed to generate data.
 Diffusion Models for Image-to-Video Generation.
-Given any image ${\bm{y}}_{0}$ from the open domain, the goal of I2V is to generate a video $X_{0}=\{{\bm{x}}_{0}^{i}\}_{i=1}^{N}$ with dynamic and natural motion while maintaining the appearance of ${\bm{y}}_{0}$.
-This task can be formulated as designing a conditional distribution $p_{\theta}(X_{0}|{\bm{y}}_{0})$, which be achieved by a conditional diffusion model:
+Given an image ${\bm{y}}_{0}$ from the open domain, the goal of I2V is to generate a video $X_{0}=\{{\bm{x}}_{0}^{i}\}_{i=1}^{N}$ with dynamic and natural motion while keeping alignment with the appearance of ${\bm{y}}_{0}$.
+This task can be formulated as designing a conditional distribution $p_{\theta}(X_{0}|{\bm{y}}_{0})$, which is achieved by a conditional diffusion model minimizing:
 $\mathbb{E}_{X_{0},{\bm{y}}_{0},{\bm{\epsilon}},t}\left[\|{\bm{\epsilon}}_{\theta}(X_{t},{\bm{y}}_{0},t)-{\bm{\epsilon}}\|_{2}^{2}\right],$ (3)
@@
 Typically, ${\bm{y}}_{0}$ is the first frame of $X_{0}$ and DynamiCrafter [63] adopts a randomly selected frame from $X_{0}$ as ${\bm{y}}_{0}$.
-The key issue is effectively integrating the conditional image ${\bm{y}}_{0}$ into the diffusion model.
+The key issue is to effectively integrate the conditional image ${\bm{y}}_{0}$ into the diffusion model.
 Most methods use CLIP image embeddings [40] to maintain the semantic content of ${\bm{y}}_{0}$.
 Notably, VideoCrafter1 [12] and Dynamicrafter [63] employ the last layer’s full patch visual tokens from the CLIP ViT, enriching the encoded information, and other approaches prefer the class token layer.
-Yet, solely depending on these embeddings, such as in VideoCrafter1 [12], compromises detail retention, impacting image alignment.
+Yet, solely depending on these embeddings, such as in VideoCrafter1 [12], compromises detail retention, resulting in degradation of the image alignment.
 To enhance detail representation, I2VGen-XL [68] combines the conditional image with the noisy initial frame, while VideoComposer [57] develops a STC-encoder for multiple conditions.
 Although superior to CLIP image embeddings, these strategies still fail to fully retain the conditional image content.
-To mitigate this, AnimateAnything [15], Dynamicrafter [54] and SVD [9] choose to directly concatenate noisy video $X_{t}$ with ${\bm{y}}_{0}$, which provides detailed information.
-Figure 2: Identifying conditional image leakage in I2V-DMs.
-We corrupt a ground truth (GT) clean video and use it to compute $\hat{X}_{t\rightarrow 0}$ prediction across time steps.
-As time progresses, $\hat{X}_{t\rightarrow 0}$ exhibits markedly reduced motion than GT, with the videos becoming nearly static at the final step $T$, suggesting the existence of conditional image leakage in a large time step.
-3 Conditional Image Leakage
-In this section, we first identify a commonly overlooked issue in I2V-DMs: conditional image leakage, discussed in Sec. 3.1. We then discuss existing I2V-DMs with respect to this issue in Sec. 3.2.
+To mitigate this, AnimateAnything [15], Dynamicrafter [63] and SVD [9] directly concatenate noisy video $X_{t}$ with ${\bm{y}}_{0}$, which injects detailed information to the model.
+Apart from the prior work mentioned before, we discuss other related work about diffusion models for image generation and video generation in the Appendix F.
+Figure 2: Identifying conditional image leakage.
+As time step progresses, the noisy input becomes heavily corrupted, whereas the conditional image retains considerable detail from GT.
+This biases the model to over-rely on the conditional image at large $t$, resulting in videos with less motion than GT.
+3 Method
+Although existing I2V-DMs discussed in Sec.
+2 have achieved significant progress, such models are not fully understood.
+In this section, we first identify a critical yet previously overlooked issue in I2V-DMs: conditional image leakage (CIL) (see Sec.
+3.1).
+We then address this issue from both inference and training aspects accordingly (see Sec.
+3.2).
+Finally, we offer insights into existing I2V-DMs through the lens of CIL (see Sec.
+3.3).
 3.1 Identifying Conditional Image Leakage in Image-to-video Diffusion Models
-Although existing I2V-DMs discussed in Sec.
-2 have obtained significant progress, such models are not fully understood.
-This section revisits and underscores a critical yet previously overlooked issue: conditional image leakage (CIL).
 As shown in Fig.
-1, ideally, I2V-DMs should predict the added noise ($\epsilon$-prediction) or clean video (${\bm{x}}_{0}$-prediction) mainly from noisy input $X_{t}$, using the conditional image ${\bm{y}}_{0}$ as auxiliary content guidance.
-However, as the diffusion process progresses—particularly at large time steps, the heavily corrupted input $X_{t}$ contains minimal information about $X_{0}$, making it challenging to predict clean frames.
-Meanwhile, the conditional image retains considerable detail of $X_{0}$.
-This biases the model to over-rely on the conditional image while neglecting the essential task of synthesizing video from noisy inputs.
-For instance, a diffusion model that takes a shortcut by directly replicating ${\bm{y}}_{0}$ can even achieve a lower loss than one that accurately predicts the clean video from a heavily corrupted $X_{t}$.
-As a result, the synthesized videos lack dynamic and vivid motion.
-Notably, recent strategies that involve adjusting the noise schedule towards more noise [14, 26, 9, 31] may further exacerbate this issue (see details in Appendix C).
-To validate this issue, we corrupt a ground truth (GT) clean video $X_{0}$ using the forward transition kernel in Eq.
+1, we observe that, regardless of the input motion scores, the motion scores of generated videos from existing I2V-DMs [9, 69] are consistently lower than expected.
+This raises the question: why do these models always produce lower motion scores, rather than fluctuating above or below the expected values, as observed in our method?
+To understand this, we need to consider the source of motion information in the generated videos, which comes from the noisy input $X_{t}$.
+Ideally, I2V-DMs should rely primarily on $X_{t}$ for motion, with the static conditional image ${\bm{y}}_{0}$ providing content guidance.
+However, as shown in Fig.
+2, at large time steps, the noisy input $X_{t}$ becomes increasingly corrupted, while the conditional image ${\bm{y}}_{0}$ retains significant information of the target video.
+This biases the model to over-rely on the conditional image and neglect the noisy inputs, leading to videos with reduced motion.
+To validate this, we corrupt a ground truth (GT) clean video $X_{0}$ via the forward transition kernel in Eq.
 (1) and use it as the noisy input to compute the one-step prediction $\hat{X}_{t\rightarrow 0}$ at time $t$:
 $\hat{X}_{t\rightarrow 0}=\left(X_{t}-\sigma_{t}{\bm{\epsilon}}_{\theta}(X_{t},y,t)\right)/\alpha_{t}.$ (4)
-Ideally, $\hat{X}_{t\rightarrow 0}$ attempts to predict the GT $X_{0}$ from noisy input $X_{t}$ and exhibit comparable motion dynamics.
+Ideally, $\hat{X}_{t\rightarrow 0}$ should predict the GT $X_{0}$ from noisy input $X_{t}$ and exhibit comparable motion dynamics.
 However, as shown in Fig.
-2, as time progresses—particularly at large time steps, $\hat{X}_{t\rightarrow 0}$ exhibits markedly reduced motion than GT, with the videos becoming nearly static at the final step $T$, suggesting the existence of the conditional image leakage.
-Consequently, it will generate an almost still video starting from $T$ by duplicating the conditional image (see Fig.
-6).
-Figure 3: Understanding exiting work from conditional image leakage.
-(a) I2VGen-XL [58] mitigates the leakage at the expense of image alignment.
-(b) Increasing the SVD motion score induces unnatural camera movements with stationary objects.
-Our inference strategy, however, produces dynamic and natural motion in both people and the teapot.
-3.2 Understanding Existing Work from Conditional Image Leakage
-We now examine three popular I2V-DMs with respect to conditional image leakage: I2VGEN-XL [58], Dynamicrafter [54], and SVD [9].
-Although these models do not explicitly address conditional image leakage, we believe their strategies may mitigate it to some extent.
-Firstly, I2VGEN-XL [58] adds ${\bm{y}}_{0}$ with the first frame of noisy input, which reduces information about ${\bm{y}}_{0}$ largely and thus mitigates the leakage issue.
-However, Fig.
-3(a) indicates that the generated videos do not accurately preserve ${\bm{y}}_{0}$.
-Dynamicrafter [54] and SVD [9] directly feed ${\bm{y}}_{0}$ into I2V-DMs, enhancing detail preservation but also facilitating leakage.
-Dynamicrafter selectively refines spatial layers while preserving the pre-trained temporal layers, which contain motion priors and thus maintain motion dynamics to a certain extent.
-However, it does not inherently solve the leakage issue.
-SVD [9] introduces an additional motion score to compel the model to generate videos with motion to align with motion score, thereby reducing reliance solely on ${\bm{y}}_{0}$ and mitigating conditional image leakage.
-However, we argue that it does not address the fundamental challenge facing I2V-DMs, where I2V-DMs should generate clean video primarily from noisy input.
-An example in Fig.
-3(b) highlights this limitation.
-By increasing the motion score, the method produces videos with significant camera movement to meet the scalar motion score but leaves objects static.
-In contrast, the strategy we propose in Sec.
-4.1 results in a video with vivid and naturalistic movement, where both the people and the teapot exhibit dynamic motion, offering a more effective solution to the fundamental challenges of I2V-DMs.
-Furthermore, if an inappropriate motion score is provided during inference, the resulting video may exhibit unnatural motion or poor temporal consistency to match the given score (see Fig.
-12 in the Appendix).
-In summary, despite efforts by these works to somewhat mitigate conditional image leakage, the issue persists, underscoring the ongoing need for further exploration.
-Figure 4: Benefits of tuning $M$ and Analytic-Init w.r.t.
-visual quality.
-An appropriate $M$ can enhance performance by increasing motion without compromising other performance.
-A too-small $M$ delivers poor visual quality due to the training-inference gap.
-Analytic-Init helps to mitigate it.
-4 Solving Conditional Image Leakage in Image-to-video Diffusion Models
-In this section, we tackle the issue of conditional image leakage from two perspectives: first, by outlining a training-free inference strategy to mitigate it, and second, by presenting a training strategy designed to rectify this phenomenon during the training phase.
-4.1 Inference Strategy
+2, as time progresses—particularly at large time steps, $\hat{X}_{t\rightarrow 0}$ exhibits markedly reduced motion than GT, indicating the conditional image leakage.
+This results in videos with reduced motion starting from time $T$.
+Notably, recent techniques that adjust the noise schedule towards higher noise levels [14, 27, 9, 32] may further exacerbate this issue (see Appendix C for details).
+Figure 3: Benefits of Analytic-Init.
+(a) An early start time $M$ enhances motion but a too-small $M$ degrades visual quality due to the training-inference gap, which Analytic-Init helps to reduce.
+(b) Analytic-Init produces higher motion scores with lower errors, mitigating conditional image leakage.
+3.2 Solving Conditional Image Leakage in Image-to-video Diffusion Models
+Building upon the above analysis, this section presents general strategies to address the issue of conditional image leakage in both the inference and training aspects.
+Inference strategy.
 As discussed in Sec.
-3.1, conditional image leakage commonly occurs at large time steps.
-To address it, a straightforward solution is to start the generation process from an earlier time step $M\in[0,T]$, avoiding the unreliable later stages of I2V-DMs.
+3.1, conditional image leakage easily occurs at large time steps.
+To this end, a straightforward solution is to start the generation process from an earlier time step $M\in(0,T)$, thus avoiding the unreliable later stages of I2V-DMs.
 Let $p_{M}(X_{M})$ denote the initial noise distribution at the start time $M$.
 Initially, we set $p_{M}(X_{M})=p_{T}(X_{T})$, i.e.
-${\mathcal{N}}({\bm{0}},{\bm{I}})$ in VP-SDE [23] or ${\mathcal{N}}({\bm{0}},\sigma_{T}^{2}{\bm{I}})$ in VE-SDE [28].
-As illustrated in the third column ($M=0.92T$) of Fig.
-4, this straightforward strategy markedly improves motion dynamics without sacrificing other performance.
-However, a smaller $M$ value (e.g., $M=0.8T$) hurts the visual quality due to the training-inference discrepancy.
-To bridge the training-inference gap, we propose to minimize the KL divergence between the initial noise distribution $p_{M}(X_{M})$ and the actual marginal distribution of the diffusion process at time $M$, denoted as $q_{M}(X_{M})$.
-Inspired by existing work [4, 3, 55], when $p_{M}(X_{M})$ is modeled as a normal distribution ${\mathcal{N}}(X_{M};{\bm{\mu}}_{p},\sigma_{p}^{2}{\bm{I}})$, we derive analytical expressions for the optimal mean ${\bm{\mu}}_{p}^{*}$ and variance $\sigma_{p}^{2\ast}$ as follows.
-We refer to it as the Analytic Noise Initialization (Analytic-Init).
+${\mathcal{N}}({\bm{0}},{\bm{I}})$ in VP-SDE [24, 63, 12, 69] or ${\mathcal{N}}({\bm{0}},\sigma_{T}^{2}{\bm{I}})$ in VE-SDE [29, 9].
+As illustrated in Fig.
+3 (a), this straightforward strategy markedly improves motion dynamics without sacrificing other performance.
+However, a smaller $M$ value (e.g., $M=0.8T$) results in poor visual quality due to the training-inference discrepancy.
+To mitigate this gap, we propose Analytic Noise Initialization (Analytic-Init) to refine the initial noise distribution $p_{M}(X_{M})$ by minimizing the KL divergence between it and the true marginal distribution $q_{M}(X_{M})$ of the forward diffusion process.
+Inspired by previous work [4, 3, 64], we demonstrate that when $p_{M}(X_{M})$ is modeled as a normal distribution ${\mathcal{N}}(X_{M};{\bm{\mu}}_{p},\sigma_{p}^{2}{\bm{I}})$, the optimal mean ${\bm{\mu}}_{p}^{*}$ and variance $\sigma_{p}^{2\ast}$ have analytical solutions, as stated in Proposition 1.
+(a) TimeNoise $p_{t}(\beta_{s})$ with hyperparameters $\beta_{m}$ (maximum noise) and $\mu(t)$ (distribution center).
+(b) Distribution center $\mu(t)=2t^{a}-1$ with $a$ controlling monotonic behavior flexibly.
+(c) Effects of tuning $a$ in $\mu(t)$ and $\beta_{m}$, with higher noise levels enhancing motion but reducing alignment.
+Figure 4: Visualization of TimeNoise and the impact of tuning its hyperparameters.
+(a) The designed $p_{t}(\beta_{s})$ favors high noise levels at large $t$, gradually shifting to lower noise levels as $t$ decreases.
+This is achieved by (b) $\mu(t)$ increasing monotonically with $t$.
+Finally, (c) modifying $a$ and $\beta_{m}$ enables a trade-off between dynamic motion and image alignment.
 Proposition 1.
@@
 where $q(X_{0})$ denotes the data distribution , $d$ denotes the dimension of the data, and $X_{0}^{(j)}$ denotes the j-th component of $X_{0}$.
-Appendix A provides detailed proof of Proposition 1.
-Empirically, we can estimate ${\bm{\mu}}_{p}^{*}$ and $\sigma_{p}^{2\ast}$ using the method of moments [4].
-Compared with the naive approach where $p_{M}(X_{M})=p_{T}(X_{T})$, Analytic-Init can achieve superior visual quality by reducing the training-inference gaps, especially with a smaller $M$.
-The theoretical insights and Proposition 1 are validated by the results in Tab.
-4, measured by the commonly used FVD and IS metrics.
-Figure 5: Effects of tuning $\mu(t)$ and $\beta_{m}$ in TimeNoise.
-Adjusting these parameters allows for a flexible trade-off between dynamic motion and image alignment.
-Higher noise levels result in more pronounced motion but reduced image alignment.
-Figure 6: Effectiveness of TimeNoise to address conditional image leakage.
-We replicate the experiments described in Sec.
-3.1 to evaluate TimeNoise.
-Applying TimeNoise ensures that $\hat{X}_{t\rightarrow 0}$ maintains motion dynamics comparable to the GT across all time steps, indicating the absence of conditional image leakage.
-Consequently, we can generate videos with high motion starting from $T$.
-4.2 Training Strategy
-The previous section explores strategies to mitigate conditional image leakage during inference.
-This section presents methods to address this issue during the training phase.
-As noted in Sec.
-3.1, the conditional image ${\bm{y}}_{0}$ remains informative of the target video, leading I2V-DMs to depend excessively on it.
-Therefore, an intuitive idea is to perturb ${\bm{y}}_{0}$ to relieve the reliance on it.
-Our first attempt is to introduce noise similar in scale to that used in $X_{t}$.
-This aims to equalize the model’s challenge in predicting clean video from $X_{t}$ or ${\bm{y}}_{t}$, reducing its dependency on ${\bm{y}}_{0}$.
-However, this strategy also makes it difficult to harness ${\bm{y}}_{0}$, resulting in lower video quality.
+The proof of Proposition 1 is provided in Appendix A.
+Empirically, ${\bm{\mu}}_{p}^{*}$ and $\sigma_{p}^{2\ast}$ can be estimated using the method of moments [4].
+The steps for inference are outlined in Algorithm 1.
+As demonstrated by the qualitative results in Fig.
+3 (a) and quantitative results in Tab.
+8, Analytic-Init improves video quality by reducing the training-inference gap, especially for smaller $M$.
+Finally, as shown in Fig.
+3 (b) and Tab.
+1, Analytic-Init produces higher motion scores with lower errors, allowing for more accurate motion control and reducing conditional image leakage.
+Figure 5: Benefits of TimeNoise.
+TimeNoise effectively reduces conditional image leakage by (a) generating $\hat{X}_{t\rightarrow 0}$ that maintains motion dynamics comparable to the GT across all time steps, and (b) achieving higher motion scores with lower errors.
+Training strategy.
+In this section, we show how to address the issue of conditional image leakage during the training phase.
+As outlined in Sec.
+3.1, the conditional image ${\bm{y}}_{0}$ retains substantial details of the target video, causing I2V-DMs to rely heavily on it.
+To mitigate this, a natural approach is to perturb ${\bm{y}}_{0}$ to relieve this dependency.
+Our first attempt is to introduce noise at a similar scale to that in $X_{t}$, aiming to balance the model’s challenge of predicting clean video from $X_{t}$ or ${\bm{y}}_{0}$, thereby lessening reliance on ${\bm{y}}_{0}$.
+However, this strategy also makes it difficult to employ ${\bm{y}}_{0}$, resulting in lower video quality.
 To overcome this, we propose a noise distribution on ${\bm{y}}_{0}$ that introduces substantial noise to prevent leakage while maintaining a cleaner ${\bm{y}}_{0}$ to aid content generation.
-As discussed earlier, as time progresses, $X_{t}$ becomes less informative about $X_{0}$, increasing the risk of leakage.
-This insight leads us to develop a time-dependent noise distribution $p_{t}(\beta_{s})$ (TimeNoise).
-The design principle is to favor high noise levels at large time steps to sufficiently disrupt ${\bm{y}}_{0}$, shifting towards lower noise levels as the time step decreases.
+Given that $X_{t}$ contains less information about $X_{0}$ as time progresses, increasing the risk of leakage, we further develop a time-dependent noise distribution $p_{t}(\beta_{s})$ (TimeNoise).
+The key principle is to favor high noise levels at large time steps to sufficiently disrupt ${\bm{y}}_{0}$, shifting towards lower noise levels as the time step decreases.
 To achieve this, we employ a logit-normal distribution [18, 1] defined as below:
@@
 where $\text{logit}(\frac{\beta_{s}}{\beta_{m}})$ follows a normal distribution centered around $\mu(t)$ with a standard deviation of 1.
-Our noise distribution includes two hyperparameters: $\beta_{m}$, the maximum noise level, and $\mu(t)$, the center of the distribution.
+This noise distribution includes two hyperparameters: $\beta_{m}$, the maximum noise level, and $\mu(t)$, the center of the distribution.
 As illustrated in Fig.
-8(a), we can adhere to the previously stated design principle of $p_{t}(\beta_{s})$ by adjusting $\mu(t)$ over time $t$.
+4 (a), we can adjust $\mu(t)$ over time $t$ to satisfy the previously mentioned design principle.
 Finally, the noisy conditional image ${\bm{y}}_{s}$ at time $t$ is obtained by ${\bm{y}}_{s}={\bm{y}}_{0}+\beta_{s}{\bm{\epsilon}}$, where $\beta_{s}\sim p_{t}(\beta_{s}),{\bm{\epsilon}}\sim{\mathcal{N}}({\bm{0}},{\bm{I}})$.
 We also tried other adding noise choices but found them to be less effective (see Appendix D).
-Next, we conduct a systematic analysis of two hyperparameters to investigate their impact on video generation.
-The first, $\mu(t)$, should ideally increase monotonically with time, where we set it ranging from $\mu(0)=-1$ to $\mu(T)=1$.
-We explore three typical functional forms for $\mu(t)$ (see Fig.
-8(b)): (1) linear: $\mu(t)=2t-1$, (2) convex:$\mu(t)=2t^{a}-1$, where $a>1$ and (3) concave: $\mu(t)=2t^{a}-1$, where $a\in(0,1)$.
+During inference, we add a fixed noise level to the conditional image across all time steps, following CDM[25], as the model is trained with varying noise levels.
+Empirically, we find directly using the clean conditional image performs well and thus adopt it for simplicity.
+The full algorithm is shown in Algorithm 2.
+Next, we conduct a systematic analysis of the two hyperparameters to investigate their impact on video generation.
+Firstly, $\mu(t)$ is designed to increase monotonically with time, ranging from $\mu(0)=-1$ to $\mu(T)=1$.
+To formalize this, we define $\mu(t)$ as a power function: $\mu(t)=2t^{a}-1$, where $a>0$.
+This formulation allows flexible tuning of $a$ to control the monotonic behavior, where smaller values of $a$ cause higher noise levels to be sampled at later time steps.
 As shown in Fig.
-5(a), our TimeNoise strategy produces videos with more dynamic motion than the baseline.
-Functions that introduce higher noise levels, such as the concave function, enhance dynamic motion but reduce temporal consistency and image alignment.
-We also test the following baselines: (1) constant, where $\beta_{s}=\beta_{m}\frac{\mu(t)+1}{2}$ represents a single value rather than a distribution, and (2) a conditioning augmentation method from CDM [24].
-Results indicate that the former method results in poor image alignment, while the latter shows low motion (see Fig.
-10 in the Appendix).
-We then examine the maximum noise level, $\beta_{m}$.
+4 (b), (1) $a=1$ corresponds to a linear increase; (2) $a\in(0,1)$ represents a concave function, indicating a faster noise level increase; and (3) $a>1$ corresponds to a convex function, indicating a slower increase in noise levels over time.
 As shown in Fig.
-5(b), a higher $\beta_{m}$ enhances dynamic motion but decreases temporal consistency and image alignment, while a lower $\beta_{m}$ leads to less motion.
-Considering that VideoCrafter1 and DynamiCrafter utilize full patch visual tokens from CLIP, not just the class token, we also apply noise to these tokens.
-Finally, we replicated the experiments outlined in Sec.
-3.1, and the results depicted in Fig.
-6 demonstrate that our TimeNoise ensures $\hat{X}_{t\rightarrow 0}$ maintains motion dynamics comparable to the ground truth across all time steps.
-This indicates an absence of conditional image leakage, underscoring the effectiveness of our approach.
-Figure 7: Visualizing the effectiveness of our inference and training strategies.
+4 (c), higher noise levels (e.g., when $a<1$) lead to increased dynamic motion but reduced temporal consistency and image alignment.
+For the maximum noise level $\beta_{m}$, the only constraint is that it must be greater than 0.
+As shown in Fig.
+4 (c), a higher $\beta_{m}$ enhances dynamic motion but decreases temporal consistency and image alignment, while a lower $\beta_{m}$ reduces motion.
+Additionally, we apply TimeNoise to the CLIP Image Embedding for both VideoCrafter1 [12] and DynamiCrafter [63], as they use full patch visual tokens from CLIP, which contain substantial information about the conditional image, increasing the likelihood of leakage.
+Finally, we replicate the experiments described in Sec.
+3.1, with results presented in Fig.
+5 and Tab.
+1.
+These results demonstrate that TimeNoise achieves higher motion scores with reduced error and ensures that $\hat{X}_{t\rightarrow 0}$ maintains motion dynamics comparable to the ground truth across all time steps, effectively mitigating conditional image leakage.
+Figure 6: Understanding exiting work from conditional image leakage.
+I2VGen-XL [68] and VideoCrafter1 [12] mitigates the leakage at the expense of image alignment.
+The SVD produces videos with camera movements while keeping objects relatively static to meet high motion scores, while ours generates videos that feature both natural and dynamic object movements.
+Figure 7: Qualitative results of TimeNoise and Analytic-Init applied to various I2V-DMs.
 Ours significantly enhances video dynamism while maintaining image alignment and temporal consistency.
@@
 denote VideoCrafter1 [12] and DynamiCrafter [63] respectively.
-5 Related Work
-Diffusion Models for Image Generation.
-Recently, diffusion models have recently achieved significant breakthroughs in image, video, and 3D generation [23, 16, 40, 37, 5, 59, 36, 52, 53, 60].
-For image generation, the latent diffusion model [40] addresses computational costs by leveraging VQ-VAE [47] to transform pixel-space images into compact latent representations, subsequently training diffusion models within this latent space.
-Building upon the latent space concept, subsequent works such as SDXL [38], DALLE-3 [8], and SD3 [18] have further enhanced performance.
-Exploiting these advancements in text-to-image diffusion models, numerous methods have demonstrated promising results in text-driven controllable image generation and image editing [57, 21, 46].
-In image editing, studies such as Prompt-to-Prompt [21] and Plug-and-Play [46] have explored attention-based control mechanisms over generated content, consistently delivering impressive results.
-For image translation, EGSDE [59] and DiffuseIT [29] propose to employ an additional energy function to guide the inference process of a pre-trained diffusion model.
-Diffusion Models for T2V Generation.
-Approaches to T2V generation can be classified into two main categories.
-The first category involves directly generating videos based on text [42, 10, 22, 12, 13, 48, 56, 50, 9, 17].
-For instance, as a pioneering endeavor, Make-A-Video [42] utilizes a pre-trained text-to-image model along with a prior network for T2V diffusion models, obviating the necessity for paired video-text datasets.
-VideoLDM [10] maintains the parameters of a pre-trained T2I model while fine-tuning newly introduced temporal layers.
-More recently, transformers have emerged as a foundational architecture for video generation due to their scalability [20, 35, 34].
-The second category typically entails a two-step generation process: first generating an image based on the textual input and subsequently creating a video conditioned on the text and the generated image [30, 19].
-For instance, Emu-video [19] initializes a factorized text-to-video model using a pre-trained text-to-image model and then fine-tunes temporal modules in the I2V stage.
+3.3 Understanding Existing Work from Conditional Image Leakage
+In this section, we analyze popular I2V-DMs [63, 9, 69, 12, 68] through the lens of CIL.
+Although these models do not explicitly address this issue, we believe their strategies mitigate it to some degree.
+Firstly, some methods only use partial information from the conditional image, which can help reduce leakage.
+For example, VideoCrafter1 [12] only utilizes CLIP Image Embedding, and I2VGEN-XL [68] adds ${\bm{y}}_{0}$ to the first frame of the noisy input.
+However, as shown in Fig.
+6, the videos generated by these methods do not fully capture the details of ${\bm{y}}_{0}$.
+To address this, models like Dynamicrafter [63] directly incorporate ${\bm{y}}_{0}$ into I2V-DMs, improving detail preservation but also increasing the risk of leakage.
+Dynamicrafter selectively refines spatial layers while preserving the pre-trained temporal layers, which contain motion priors and thus maintain motion dynamics to a certain extent.
+However, it does not inherently solve the leakage issue.
+Moreover, some methods introduce external signals [9, 69, 65, 62, 60], forcing the model to align with additional conditions, which reduces its dependence on ${\bm{y}}_{0}$ and helps mitigate leakage.
+However, we argue that they do not address the core challenge of I2V-DMs, which should predict clean video primarily from noisy input to capture motion information.
+As illustrated in Fig.
+6, the SVD often results in static objects with excessive camera movements to meet high motion score requirements.
+In contrast, our method generates videos with natural, vivid object movements.
+In summary, while the above methods mitigate CIL to some extent, our approach provides a more effective solution to the fundamental challenges in I2V-DMs, enabling more precise motion control and enhancing naturalness by focusing more on the noisy input.
+Table 1: Quantitative results of TimeNoise and Analytic-Init applied to various I2V-DMs on the UCF101 dataset.
+<Method>-naive-tune represents a naively fine-tuned baseline using the same training setup as our TimeNoise, ensuring a fair comparison.
+<Method>-CIL denotes the full version using both TimeNoise and Analytic-Init.
 […]
-[…]
-Figure 8: (a)Visualization of designed noise distribution.
-Adjusting $\mu(t)$ over time $t$ can satisfy the design principle.
-For example, set large $\mu(t)$ for large $t$ (purple curve) and small $\mu(t)$ for small $t$ (blue curve).
-(b)Visualization of different $\mu(t)$.
+DymiCrafter [13] 363.8 16.39 50.96
+DymiCrafter + Analytic-Init 316.3 17.66 71.04
+DymiCrafter-naive-tune 382.5 21.12 31.68
+DymiCrafter-naive-tune + Analytic-Init 342.9 22.71 50.08
+DymiCrafter-TimeNoise 334.9 21.42 72.32
+DymiCrafter-CIL 332.1 22.84 73.92
+VideoCrafter1 [12] 353.9 18.75 63.36
+VideoCrafter1 + Analytic-Init 341.6 19.86 139.04
+VideoCrafter1-naive-tune 460.3 23.98 62.72
+VideoCrafter1-naive-tune + Analytic-Init 450.1 24.50 65.12
+VideoCrafter1-TimeNoise 452.2 24.62 64.80
+VideoCrafter1-CIL 443.7 25.11 66.7
+SVD [9] 388.3 36.32 16.64
+SVD + Analytic-Init 382.0 36.81 19.68
+SVD-naive-tune 311.0 22.03 9.60
+SVD-naive-tune + Analytic-Init 277.1 22.18 20.64
+SVD-TimeNoise 272.2 23.01 20.96
+SVD-CIL 272.4 25.18 21.44
 4 Experiments
-Datasets and Evaluation Metrics.
-We validate our training strategy on the WebVid-2M dataset [2], resizing and center-cropping all videos to $320\times 512$ with 16 frames.
-For evaluation, we validate our methods on UCF101 and our collected ImageBench, comprising 100 images from various websites and T2I models like SDXL [38] and UniDiffuser [5], covering categories like nature, humans, and more for comprehensive evaluation.
+4.1 Setup
+Datasets.
+We use WebVid-2M [2] as the training dataset, with all videos resized and center-cropped to $320\times 512$ at 16 frames and 3 fps.
+For evaluation, we use UCF101 [49] and our ImageBench dataset, which includes diverse categories (e.g., nature, humans, animals, plants, food, vehicles) and complex elements like numerals, colors, and intricate scenes, similar to DrawBench [44].
+In total, 100 images are collected from various websites and T2I models such as SDXL [39] and UniDiffuser [5].
+Evaluation Metrics.
+On UCF101, we report Fréchet Video Distance (FVD), Inception Score (IS), and Motion Score (MS).
 For ImageBench, we conduct user studies with 10 subjects to perform pairwise comparisons of our methods against baselines, evaluating motion, temporal consistency, image alignment, and overall performance.
-On UCF101, we report Fréchet Video Distance (FVD) and Inception Score (IS).
-Further details can be found in Appendix B.
+For motion-conditioned methods, we also report the motion score error between the generated video and the input motion score at various levels.
+Motion scores are computed using flow maps following SVD [9], except for the PIA [69], where we follow its original algorithm and compute the $L1$ distance.
+More details on the metric computations are provided in Appendix B.
 Implementation Details.
@@
 The function $\mu(t)=2t^{5}-1$ is applied across all baselines.
-These baselines are fine-tuned for 20,000 iterations, using either the official code or a replication of the official settings except for batch size.
+These baselines are fine-tuned for 20,000 iterations, using either the official code or replication of the official settings except for batch size.
 More detailed information can be found in the Appendix B.
-[…]
+Figure 8: User preference percentages for ours against baselines.
+<Method>-NT.
+is the naively fine-tuned baseline.
+DC.
+and VC.
+denote DynamiCrafter and VideoCrafter1.
+Our strategies significantly enhance video dynamism while maintaining image alignment and temporal consistency, thus achieving superior results overall.
 4.2 Results
 The effectiveness of our inference and training strategy.
-We validate our strategies on three I2V-DMs: VideoCrafter1 [12], DynamiCrafter [54], and SVD [48].
-Demonstrated in Fig.
-7, our strategies significantly enhance video dynamism and natural motion, all while maintaining image alignment and temporal consistency.
-The findings are corroborated by the user study in Fig.
-9.
-Considering overall aspects, ours achieves superior results.
-Tab.
-8 showcases the quantitative superiority of our inference and training strategies, which outperform all baselines significantly.
-For instance, on DynamiCrafter [54], we improve the FVD score by 47.5 and 47.6 using our inference and training strategy respectively, suggesting the effectiveness of our method.
+We validate our strategies on following I2V-DMs: VideoCrafter1 [12], DynamiCrafter [63] and SVD [54].
+We validate our inference strategy against the original method and a naively finetuned version, and our training strategy against a naively finetuned version.
+The quantitative comparisons and qualitative results are presented in Tab.
+1, Fig.
+3, Fig.
+5 and Fig.
+7, leading to several key observations.
+First, our strategies significantly improve motion scores and reduce motion score error, demonstrating their effectiveness in precisely controlling motion degrees and mitigating conditional image leakage.
+Second, we achieve both dynamic motion and high video quality, as reflected by improved FVD and IS scores.
+The user study in Fig.
+8 further supports that our strategies enhance video dynamism while preserving image alignment and temporal consistency, yielding superior overall performance.
 Evaluating the combined inference and training strategies.
@@
 As illustrated in Tab.
-3, for DynamiCrafter [54] and SVD [9], our TimeNoise effectively mitigates conditional image leakage, rendering the additional inference strategy less impactful.
+7, for DynamiCrafter [63] and SVD [9], our TimeNoise effectively mitigates conditional image leakage, rendering the additional inference strategy less impactful.
 Conversely, for VideoCrafter1 [12], which relies solely on CLIP image embedding for information, employing excessive TimeNoise disrupts image alignment.
 Hence, we utilize a moderate TimeNoise ($\beta_{m}=25$), where the inference strategy remains effective.
+Comparison with SOTA noise initialization and conditioning augmentation methods.
+We compare our Analytic-Init with SOTA noise initialization methods [41, 61, 19] and our TimeNoise with SOTA conditioning augmentation [25].
+As shown in Tab.
+2, Analytic-Init outperforms all baselines.
+It achieves higher motion scores than FrameInit [41], which limits motion by duplicating the conditional image as static guidance.
+Compared to FreeInit [61], our method reduces inference time by more than half while slightly improving performance, as FreeInit’s iterative process increases time.
+Compared to Progressive Noise [19], Analytic-Init performs better by generating videos from an earlier time step, a gap the baseline cannot handle effectively.
+As shown in Fig.
+9 in the Appendix, CDM [25] results in lower motion, whereas our method produces more dynamic videos.
+Additionally, a naive baseline using a single value for $\beta_{s}=\beta_{m}\frac{\mu(t)+1}{2}$ leads to poor image alignment, further demonstrating the effectiveness of our time-dependent noise distribution.
+Table 2: Comparison with SOTA noise initialization methods. Rate. is the average user ranking of each method based on overall performance. Time measures the duration needed to generate a video.
+[…]
+FrameInit [41] 380.7 20.09 32.16 4.57 24.3s/it
+FreeInit [61] 347.4 22.66 46.24 1.95 49.6s/it
+Progressive Noise [19] 358.1 21.35 49.76 3.31 23.3s/it
+Analytic-Init 342.9 22.71 50.08 1.77 22.6s/it
 5 Conclusions and Discussions
@@
 According to Lemma 2 of Analytic-DPM [4], the KL divergence between initial noise distribution $p_{M}(X_{M})$ and the actual marginal distribution $q_{M}(X_{M})$ can be expressed as:
-$D_{KL}(q_{M}(X_{M})||p_{M}(X_{M}))=D_{KL}({\mathcal{N}}(X_{M};{\bm{\mu}}_{q},\Sigma_{q})||p_{M}(X_{M}))$
+$D_{KL}(q_{M}(X_{M})||p_{M}(X_{M}))=D_{KL}({\mathcal{N}}(X_{M};{\bm{\mu}}_{q},\Sigma_{q})||{\mathcal{N}}(X_{M};\mu_{p},\sigma_{p}^{2}I))$
 $+H({\mathcal{N}}(X_{M};{\bm{\mu}}_{q},\Sigma_{q}))-H(q_{M}(X_{M})),$ (7)
 where ${\bm{\mu}}_{q},\Sigma_{q}$ denote expectation and covariance matrix of $q_{M}(X_{M})$, and $H(\cdot)$ denotes the entropy of a distribution.
-Since $q_{M}(X_{M})$ is independent of $p_{M}(X_{M})$, the last two terms can be considered as constants.
-Therefore, minimizing the Eq.
-(A) is equivalent to minimizing $D_{KL}(N(X_{M};{\bm{\mu}}_{q},\Sigma_{q})||p_{M}(X_{M}))$.
-For normal distributions, $D_{KL}(N(X_{M};{\bm{\mu}}_{q},\Sigma_{q})||p_{M}(X_{M}))$ could be expanded as:
-$\frac{1}{2}[({\bm{\mu}}_{q}-{\bm{\mu}}_{p})^{T}\Sigma_{p}^{-1}({\bm{\mu}}_{q}-{\bm{\mu}}_{p})-\log\det(\Sigma_{p}^{-1}\Sigma_{q})+tr(\Sigma_{p}^{-1}\Sigma_{q})-d],$ (8)
-$=\frac{1}{2}[\frac{1}{\sigma_{p}^{2}}\lVert{\bm{\mu}}_{p}-{\bm{\mu}}_{q}\rVert^{2}+d\log(\sigma_{p}^{2})+\frac{1}{\sigma_{p}^{2}}tr(\Sigma_{q})-\log\det(\Sigma_{q})-d],$ (9)
+Since $q_{M}(X_{M})$ is determined by the forward diffusion process and is independent of the $\mu_{p}$ and $\sigma_{p}^{2}$ in $p_{M}(X_{M})$, the last two terms of Eq.
+(A) can be considered as constants.
+Therefore, the optimization objective can be formulated as:
+$\min_{\mu_{p},\sigma_{p}^{2}}D_{KL}(q_{M}(X_{M})||p_{M}(X_{M}))\Leftrightarrow\min_{\mu_{p},\sigma_{p}^{2}}D_{KL}({\mathcal{N}}(X_{M};\mu_{q},\Sigma_{q})||{\mathcal{N}}(X_{M};\mu_{p},\sigma_{p}^{2}I)).$ (8)
+According to the property of the KL divergence between two normal distributions, $D_{KL}({\mathcal{N}}(X_{M};{\bm{\mu}}_{q},\Sigma_{q})||{\mathcal{N}}(X_{M};\mu_{p},\sigma_{p}^{2}I))$ could be expanded as:
+$\frac{1}{2}[\frac{1}{\sigma_{p}^{2}}\lVert{\bm{\mu}}_{p}-{\bm{\mu}}_{q}\rVert^{2}+d\log(\sigma_{p}^{2})+\frac{1}{\sigma_{p}^{2}}tr(\Sigma_{q})-\log\det(\Sigma_{q})-d],$ (9)
 where $d$ denotes the dimension of the data. Minimizing Eq. (9) yields:
 ${\bm{\mu}}_{p}^{*}={\bm{\mu}}_{q}=\mathbb{E}_{q_{M}(X_{M})}[X_{M}].$ (10)
-Taking the derivative of Eq. 9 with respect to $\sigma_{p}^{2}$, we can obtain
-$\sigma_{p}^{2\ast}$ $=\frac{trCov_{q_{M}(X_{M})[X_{M}]+\lVert{\bm{\mu}}_{p}-{\bm{\mu}}_{q}\rVert^{2}}}{d},$ (11)
-$=\frac{trCov_{q_{M}(X_{M})[X_{M}]}}{d},$ (12)
+Taking the derivative of Eq. 9 with respect to $\sigma_{p}^{2}$ , the optimal $\sigma_{p}^{2}$ comes to:
+$\sigma_{p}^{2\ast}$ $=\frac{tr(Cov_{q_{M}(X_{M})}[X_{M}])+\lVert{\bm{\mu}}_{p}-{\bm{\mu}}_{q}\rVert^{2}}{d},$ (11)
+$=\frac{tr(Cov_{q_{M}(X_{M})}[X_{M}])}{d},$ (12)
 $=\frac{\sum_{j=1}^{d}(Var(X_{M}^{(j)}))}{d}.$ (13)
-Taking the $X_{M}$ defined in Eq. (1) into Eq. (10), the optimal ${\bm{\mu}}_{p}^{*}$ could be further represented as
+We now represent ${\bm{\mu}}_{p}^{*}$ and $\sigma_{p}^{2\ast}$ with $X_{0}$. Taking the $X_{M}$ defined in Eq. (1) into Eq. (10), the optimal ${\bm{\mu}}_{p}^{*}$ could be further represented as
 $\mu_{p}^{*}$ $=\alpha_{M}\mathbb{E}_{q(X_{0})}[X_{0}]+\mathbb{E}_{q({\bm{\epsilon}})}[{\bm{\epsilon}}],$ (14)
@@
 $Var(X_{M}^{(j)})=Var(\alpha_{M}X_{0}^{(j)}+\sigma_{M}{\bm{\epsilon}}^{(j)})$ (16)
-Given that $X_{0}$ and ${\bm{\epsilon}}$ are independent of each other, the variance $Var(X_{M}^{(j)})$ can be decomposed further:
+Given that $X_{0}$ and ${\bm{\epsilon}}$ are independent of each other, the variance $Var(X_{M}^{(j)})$ can be further decomposed as:
 $Var(X_{M}^{(j)})$ $=Var(\alpha_{M}X_{0}^{(j)})+Var(\sigma_{M}{\bm{\epsilon}}^{(j)}),$ (17)
@@
 All used codes in this paper and their licenses are listed in Tab.
-1.
+3.
 B.2 Training and Inference Details
 We utilize the official training code of DynamiCrafter (refer to Tab.
-1) to fine-tune both DynamiCrafter [54] and VideoCrafter1 [12], and reproduce the training code for SVD by ourselves.
+3) to fine-tune both DynamiCrafter [63] and VideoCrafter1 [12], and reproduce the training code for SVD by ourselves.
 Throughout the training phase, we maintained consistent settings across all models, with the sole exception of incorporating our TimeNoise component for a fair comparison.
@@
 Additional details can be found in Tab.
-B.2 and Tab.
-B.2.
+4 and Tab.
+5.
 Our experiments were conducted using A800-80G GPUs, and the computational costs are detailed in Tab.
-2.
+6.
 For inference, the official model codes were used for sampling (see Tab.
-1).
+3).
 Specifically, we employed a DDIM sampler with 50 steps for DynamiCrafter [63] and VideoCrafter1 [12], and Heun’s 2nd order method with 25 steps for SVD [9].
-[…]
-tableTraining settings for DynamiCrafter [54] and VideoCrafter1 [12].
-Config Value Optimizer AdamW Learning rate 1e-5 Weight decay 1e-2 Optimizer momentum $\beta_{1},\beta_{2}$=0.9, 0.999 Batch size 64 Training iterations 20,000
-[…]
-tableTraining settings for SVD [9]. Config Value Optimizer AdamW Learning rate 3e-5 Weight decay 1e-2 Optimizer momentum $\beta_{1},\beta_{2}$=0.9,0.999 Batch size 48 Training iterations 20,000
+Table 4: Training settings for DynamiCrafter [63] and VideoCrafter1 [12].
+Config Value
+Optimizer AdamW
+Learning rate 1e-5
+Weight decay 1e-2
+Optimizer momentum $\beta_{1},\beta_{2}$=0.9, 0.999
+Batch size 64
+Training iterations 20,000
+Table 5: Training settings for SVD [9].
+Config Value
+Optimizer AdamW
+Learning rate 3e-5
+Weight decay 1e-2
+Optimizer momentum $\beta_{1},\beta_{2}$=0.9,0.999
+Batch size 48
+Training iterations 20,000
 Table 6: Compute resources.
@@
 For DynamiCrafter [63] and VideoCrafter1 [12], which utilize text as an additional condition, we employ categories as the textual input.
+Motion Score.
+The Motion Scores of DynamiCrafter [63], VideoCrafter1 [12] and SVD are implemented following SVD [9].Specifically, we compute the motion score by resizing the video to $800\times 450$, extracting dense optical flow maps using RAFT [50] between adjacent frames, calculating the magnitude of the 2D vector for each pixel, averaging these magnitudes spatially, and then summing them across all frames.
+As for PIA [69], we calculate the motion score by making a slight modification to the affinity score proposed in  [69], namely: motion score = 1 - affinity score.
+More specifically, for each video, we calculate the L1 distance between each frame $v^{i}$ and the condition frame $v^{1}$ in HSV space, which is denoted as $d^{i}$.
+Then we apply this operation to all frames of video clips in the dataset and find the maximum distance value $d_{max}$.
+We normalize the distance $d^{i}$ to $[0,1]$ via $d_{max}$.
+Finally, the motion score for each frame can be calculated by $m^{i}=d^{i}/d_{max}\times(m_{max}-m_{min})$, where $m_{max}$ and $m_{min}$ are hyperparameters set as $1,0.2$ respectively.
 User study.
@@
 Surprisingly, as shown in Fig.
-11, we observe that larger $P_{mean}$ values correspond to reduced motion in the generated videos.
+10, we observe that larger $P_{mean}$ values correspond to reduced motion in the generated videos.
 We hypothesize that this is due to the increased noise added to $X_{t}$ by larger $P_{mean}$, making it more challenging to predict clean frames and thus more prone to conditional image leakage.
@@
 Figure 9: The qualitative comparison between our TimeNoise and baselines mentioned in Sec.
-4.2.
+3.2.
 The constant results in poor image alignment, while the CDM [25] shows low motion.
@@
 50 295.5 21.89
+Appendix E Algorithms for Analytic-Init and TimeNoise
+Algorithm 1 Sampling from an I2V diffusion model with Analytic-Init
+0:  the conditional image ${\bm{y}}_{0}$, the sampler for diffusion models $Sampler(\cdot,\cdot,\cdot)$
+  select the initial time step $M\in(0,T)$
+  calculate ${\bm{\mu}}_{p}^{*}$ and $\sigma_{p}^{2\ast}$ in initial noise distribution according to Eq. (5)
+  $X_{M}\sim{\mathcal{N}}(X_{M};{\bm{\mu}}_{p}^{*},\sigma_{p}^{2\ast}{\bm{I}})$
+  for $t=M,...,1$ do
+     $X_{t-1}=Sampler(X_{t},{\bm{y}}_{0},t)$
+  end for
+  return  $X_{0}$
+Algorithm 2 Training an I2V diffusion model with TimeNoise
+0:  the clean video $X_{0}$, the conditional image $y_{0}$, the noise schedule $\alpha_{t},\sigma_{t}$, the time-dependent noise distribution $p_{t}(\beta_{s})$
+  repeat
+     $t\sim{\mathcal{U}}([0,T])$
+     $\beta_{s}\sim p_{t}(\beta_{s})$blacktriangleright Sampling the noise level for the conditional image
+     ${\bm{\epsilon}}_{y}\sim{\mathcal{N}}({\bm{0}},{\bm{I}})$
+     ${\bm{y}}_{s}={\bm{y}}_{0}+\beta_{s}{\bm{\epsilon}}_{y}$blacktriangleright Adding noise to the conditional image
+     ${\bm{\epsilon}}\sim{\mathcal{N}}(0,{\bm{I}})$
+     $X_{t}=\alpha_{t}X_{0}+\sigma_{t}{\bm{\epsilon}}$
+     $\theta\leftarrow\theta-\eta\nabla_{\theta}\|{\bm{\epsilon}}_{\theta}(X_{t},{\bm{y}}_{s},t)-{\bm{\epsilon}}\|_{2}^{2}$
+  until converged
+Appendix F More Related Work on Diffusion Models for Image and Video Generation
+Diffusion Models for Image Generation.
+Recently, diffusion models have achieved significant breakthroughs in image, video, and 3D generation [24, 16, 42, 38, 5, 70, 37, 58, 59, 71].
+For image generation, the latent diffusion model [42] addresses computational costs by leveraging VQ-VAE [53] to transform pixel-space images into compact latent representations, subsequently training diffusion models within this latent space.
+Building upon the latent space concept, subsequent works such as SDXL [39], DALLE-3 [8], and SD3 [18] have further enhanced the performance.
+By exploiting such advancements in text-to-image diffusion models, numerous methods have demonstrated promising results in text-driven controllable image generation and image editing [67, 22, 52].
+As for image editing, studies such as Prompt-to-Prompt [22] and Plug-and-Play [52] have explored attention-based control mechanisms over generated content, consistently delivering impressive results.
+For image translation, EGSDE [70] and DiffuseIT [30] propose to employ an additional energy function to guide the inference process of a pre-trained diffusion model.
+Diffusion Models for T2V Generation.
+Approaches to T2V generation can be classified into two main categories.
+The first one involves directly generating videos based on text [46, 10, 23, 12, 13, 54, 66, 56, 9, 17].
+For instance, as a pioneering endeavor, Make-A-Video [46] utilizes a pre-trained text-to-image model along with a prior network for T2V diffusion models, obviating the necessity for paired video-text datasets.
+VideoLDM [10] maintains the parameters of a pre-trained T2I model while fine-tuning the additionally introduced temporal layers.
+While many models are based on a U-Net [43] architecture, more recently, transformers have emerged as a foundational architecture for video generation due to their scalability [21, 36, 35].
+The second category typically entails a two-step generation process: first generating an image based on the textual input and subsequently creating a video conditioned on the text and the generated image [31, 20].
+For instance, Emu-video [20] initializes a factorized text-to-video model using a pre-trained text-to-image model and then fine-tunes temporal modules in the I2V stage.